### v1

# EtCon: Edit-then-Consolidate for Reliable Knowledge Editing

Abstract
Knowledge editing aims to update specific facts in large language models (LLMs) without full retraining. Prior efforts sought to tune the knowledge layers of LLMs, proving effective for making selective edits. However, a significant gap exists between their performance in controlled, teacher-forcing evaluations and their real-world effectiveness in lifelong learning scenarios, which greatly limits their practical applicability. This work’s empirical analysis reveals two recurring issues associated with this gap: (1) Most traditional methods lead the edited model to overfit to the new fact, thereby degrading pre-trained capabilities; (2) There is a critical absence of a knowledge consolidation stage, leaving new facts insufficiently integrated into LLMs’ inference-time behavior under autoregressive generation, thereby leading to a mismatch between parametric knowledge and actual generation behavior. To this end, we propose Edit-then-Consolidate, a novel knowledge editing paradigm that aims to bridge the gap between theoretical knowledge editing methods and their real-world applicability. Specifically, (1) our framework mitigates overfitting via Targeted Proximal Supervised Fine-Tuning (TPSFT) that localizes the edit via a trust-region objective to limit policy drift; (2) Then, a consolidation stage using Group Relative Policy Optimization (GRPO) aligns the edited knowledge with CoT-based inference policy by optimizing trajectory-level behavior under comprehensive reward signals. Extensive experiments demonstrate our framework consistently improves editing reliability and generalization under real-world evaluations, while better preserving locality and pre-trained capabilities.
1 Introduction
Large language models (LLMs) have demonstrated unprecedented capabilities across numerous tasks Guo et al. (2025); Wang et al. (2025b), serving as foundational reasoning engines for information retrieval Yang et al. (2025a), task automation agents He et al. (2025); Liu et al. (2025b); Wang et al. (2025a), and scientific research Rosen et al. (2025); Shmatko et al. (2025). However, as the external world continuously evolves, the static nature of LLMs’ pre-trained knowledge renders specific versions rapidly obsolete Zheng et al. (2025a). While retraining a large-scale LLM with updated knowledge could address this limitation, it requires substantial computational resources and pre-training data to maintain both knowledge update efficacy and general capabilities, making frequent knowledge updates impractical Mitchell et al. (2022). Knowledge editing methods Zhang et al. (2025); Scialanga et al. (2025); Li et al. (2025b); Rozner et al. (2024) have thus garnered significant attention as techniques that achieve targeted knowledge updates through localized parameter modifications while avoiding extensive resource consumption.
Knowledge editing methods can be categorized into three main paradigms: (1) Parametric in-place editing methods Meng et al. (2022a); Dai et al. (2025); Han et al. (2024); Fu et al. (2025), which directly compute weight updates and apply them to the LLM’s weight matrices, encompassing approaches such as locate-then-edit, parameter-efficient fine-tuning, and model merging—all of which preserve model architecture without requiring additional modules; (2) Meta-learning-based methods Hartvigsen et al. (2023); Li et al. (2025b); Tan et al. (2023) that train auxiliary hypernetworks to predict weight updates for specific parameters to achieve knowledge editing objectives; (3) Memory-based methods Wang et al. (2024d; a) that store new knowledge in external modules and train LLMs to activate these modules during inference involving updated knowledge. While these methods demonstrate promise in constrained evaluation scenarios such as single editing and teacher-forcing evaluations, a significant performance gap emerges in more realistic auto-regressive evaluation and lifelong editing Jiang et al. (2024); Tan et al. (2023); Chen et al. (2024), leading recent research Gu et al. (2024a); Huang et al. (2024) to question the reliability and practical utility of existing knowledge editing methods.
In this work, we conduct a comprehensive investigation into the root causes of this performance gap, focusing on Parametric In-Place Editing methods due to their high potential for practical application in lifelong learning scenarios. Through comprehensive empirical analysis, we identify two critical issues at the root of this gap. First, most of traditional methods cause edited models to overfit to newly introduced facts. This overfitting leads to excessive specialization of model parameters to editing examples, thereby degrading pre-trained general capabilities including robust reasoning, linguistic fluency, and robustness. Second, and more critically, a fundamental absence of a dedicated knowledge consolidation phase is observed. This omission results in new information being superficially encoded at the parametric level, failing to establish deep integration with the LLM’s inference-time behavior under autoregressive generation. This discordance manifests as a critical decoupling between knowledge representation and inferential activation. While the model successfully incorporates updated knowledge parametrically, it consistently fails to retrieve, activate, or apply this knowledge during autoregressive generation.
Figure 1 illustrates the core challenge we address. When existing methods edit a fact (e.g., updating Michael Jordan’s nationality), the model may parametrically encode the new information but fail to consistently apply it during reasoning. This manifests as contradictory outputs where the model simultaneously acknowledges both old and new facts, revealing a fundamental misalignment between parametric knowledge and actual generation behavior. Our Edit-then-Consolidate framework resolves this by introducing a crucial consolidation stage that aligns the edited knowledge with the model’s inference-time policy. To address the limitation illustrated above, we propose Edit-then-Consolidate (EtCon), a two-stage knowledge-editing paradigm. In the first stage, we employ Targeted Proximal Supervised Fine-Tuning (TPSFT)—a refined variant of PSFT Zhu et al. (2025) that selectively updates only the FFN layers identified as knowledge repositories. This targeted approach, combined with trust-region constraints, ensures localized edits that preserve the model’s broader capabilities. In the second stage, we introduce a critical consolidation phase using Group Relative Policy Optimization (GRPO) to align the parametric knowledge with the model’s CoT-based inference policy through trajectory-level optimization under comprehensive reward signals.
We conduct extensive experiments on three datasets with Llama-3-8B-Instruct and Qwen2.5-7B-Instruct. Under auto-regressive generation with natural stopping and an LLM-as-a-judge protocol using GPT-4.1 (OpenAI (2025)), Edit-then-Consolidate improves editing reliability and generalization by 35%-50% over strong baselines. It also significantly enhances locality while preserving critical pre-trained capabilities. Our contributions can be summarized as follows: (1) We identify that the absence of a knowledge-consolidation stage creates a critical knowledge-behavior misalignment, serving as the key bottleneck to the real-world applicability of knowledge-editing methods. (2) We propose Edit-then-Consolidate (EtCon): TPSFT for localized parametric edits, followed by GRPO for trajectory-level consolidation that aligns parametric knowledge with actual generation behavior. (3) Extensive experiments demonstrate that EtCon improves editing reliability and generalization by 40%–50%, strengthens locality, and preserves pre-trained capabilities under realistic evaluation settings.
2 Related Work
2.1 Overview of Knowledge editing methods
From the perspective of model architecture, the three paradigms introduced above can also be coarsely grouped into two families. Parametric in-place editing methods preserve the vanilla LLM architecture. The locate-then-edit paradigm Meng et al. (2022a); Dai et al. (2025); Li et al. (2024); Zhong et al. (2025); Zhang et al. (2024c) identifies knowledge locations within LLMs and modifies targeted parameters through gradient-based or analytical solutions. PEFT methods Zhu et al. (2020); Han et al. (2024); Wang et al. (2024c); Gupta et al. (2025) directly update model parameters via regularized gradient descent to achieve knowledge updates while constraining side effects Liu et al. (2025a). These approaches seamlessly integrate with existing deployment infrastructure without additional inference latency. External-assisted editing methods rely on auxiliary modules for knowledge modification. Meta-learning approaches Tan et al. (2023); Hartvigsen et al. (2023); Li et al. (2025b) train hypernetworks to generate parameter updates, while memory-based methods Hartvigsen et al. (2023); Zhang et al. (2024b); Chen et al. (2024) encode knowledge in external modules that the LLM retrieves during inference. Despite their superior performance in balancing reliability and locality, external methods introduce deployment complexity. Given these trade-offs, our work advances parametric in-place editing for lifelong knowledge editing scenarios.
2.2 Evaluation of Knowledge editing methods
Existing research Fang et al. (2024); Qi et al. (2025); Scialanga et al. (2025) predominantly evaluates the effectiveness of knowledge editing methods using a standard set of metrics. Reliability assesses the success rate of editing by calculating the percentage where . Generalization evaluates the model’s ability to generalize to new knowledge post-editing, measured by the percentage where when presented with rephrased queries pertaining to the new knowledge. Locality measures the extent to which editing a specific fact preserves the model’s responses to questions related to neighboring, unedited facts. Conventionally, the evaluation input typically consists of simple queries with identical prompt formats, without additional contextual information. For the output, edited models’ responses are often truncated to a target answer length or constrained by examples to match a specific target answer format. In generation, teacher forcing is frequently employed, feeding ground truth tokens as input during the decoding process. Recent studies, however, have highlighted the fragility of such evaluation paradigms. Consequently, this paper adopts a realistic evaluation approach for knowledge editing methods. Details of the real-world evaluation framework are in appendix A.7
3 The Missing Consolidation Stage in Knowledge Editing
Recent studies have revealed a critical performance gap in knowledge editing methods: while achieving high success rates under controlled teacher-forcing evaluation, these methods exhibit catastrophic failures in realistic auto-regressive settings. This stark discrepancy undermines their practical utility and raises fundamental questions about the effectiveness of current approaches. Through systematic investigation, we identify that this failure stems not from the editing mechanism itself, but from a fundamental architectural omission—the absence of a knowledge consolidation stage. We hypothesize that successful knowledge updating requires a two-stage process: (1) an initial parametric edit that injects new information into LLMs’ weights, followed by (2) a consolidation phase that integrates this knowledge into the LLMs’ inference-time behavior under autoregressive generation. Without consolidation, edited knowledge remains superficially encoded at the parametric level, failing to propagate to the model’s actual generation behavior.
| Method | Reli. | Gener. | Local. | |
Llama-3-8b-Instruct |
Pre-Edit | 2.8 | 2.4 | 38.6 |
| Pre-Edit(+GRPO) | 5.2 | 4.7 | 38.4 | |
| FT-M | 16.6 | 15.5 | 29.3 | |
| FT-M(+GRPO) | 62.9 | 52.7 | 24.9 | |
| ALPHAEDIT | 18.7 | 14.0 | 6.3 | |
| ALPHAEDIT(+GRPO) | 50.4 | 38.7 | 5.4 |
To validate this hypothesis, we conduct controlled experiments augmenting existing editing methods with a consolidation mechanism. Table 1 presents compelling evidence: introducing Group Relative Policy Optimization (GRPO) as a post-editing consolidation step dramatically transforms performance. For FT-M, reliability surges from 16.6% to 62.9% on Llama3-8b-Instruct, while ALPHAEDIT improves from 18.7% to 50.4%. Crucially, these gains extend to generalization metrics, indicating that consolidation enables genuine knowledge integration rather than superficial memorization. The reward trajectories in Fig. 2 further illuminate the consolidation dynamics. The monotonic increase demonstrates stable knowledge integration, where the model progressively aligns its reasoning behavior with the edited knowledge. Notably, applying GRPO directly to unedited models yields minimal improvements (Pre-Edit: 2.8% → 5.2%), confirming that consolidation requires prior parametric editing as a foundation. These findings establish a critical insight: the limitations of current knowledge editing methods arise from treating editing as a single-stage process. The Edit-then-Consolidate paradigm we propose addresses this fundamental gap, recognizing that parametric updates and behavioral alignment are complementary but distinct requirements for successful knowledge editing.
4 The Edit-then-Consolidate Framework
Building on the observational evidence in the preceding section, we posit that the limitations of current LLM knowledge-editing methods arise primarily from the lack of a principled consolidation stage that integrates edited knowledge with the model’s reasoning behavior; moreover, repeated overfitting edits can erode general abilities. To address this, we introduce Edit-then-Consolidate paradigm: Stage I employs Targeted Proximal Supervised Fine-Tuning (TPSFT) to perform localized knowledge editing under trust-region–style constraints, thereby limiting spillover while preserving pre-trained abilities; Stage II applies Group Relative Policy Optimization (GRPO) with a task-appropriate comprehensive reward to consolidate at the trajectory level under real-world evaluation signals. The remainder of this section presents the design rationale and the interaction between these two stages.
4.1 Knowledge Editing via Targeted Proximal Fine-Tuning
In this section, we introduce Targeted Proximal Supervised Fine-Tuning (TPSFT) as a refined knowledge-editing method that addresses the trilemma of reliability, locality, and generality. This approach differs from raw PSFT that update the whole LLMs, by selectively update only the FFNs of LLMs. This targeted update strategy effectively injects new knowledge while minimizing disruption to the model’s overall architecture and pre-trained capabilities. We consider a knowledge editing dataset that contains editing instances , where each context contains a question about the new fact, and is the corresponding ground-truth answer. At the start of the editing process, we have a vanilla LLM parameterized by . We partition the model’s parameters into two disjoint sets: the target FFN parameters to be edited, , and the remaining frozen parameters, , such that . The objective of TPSFT is to learn a new set of FFN parameters, , yielding an updated model where . This model must accurately generate the target answer for a given context , while minimizing disruptions to its performance on unrelated inputs. A critical innovation in our TPSFT is the use of CoT-augmented training labels. For each editing instance , we: (1) prompt the vanilla LLM to generate a CoT reasoning path for using an instruction template (see Appendix), and (2) replace the generated answer with the target new fact , yielding the training label . This design enables learning smoothed distributions over reasoning paths rather than sharp one-hot targets. More importantly, it preserves the model’s natural reasoning patterns—the model learns to reach new answers through its inherent reasoning style rather than abandoning pre-trained capabilities. This significantly reduces disruption while ensuring accurate knowledge updates.
To achieve this, we update the targeted FFN parameters by minimizing the following Targeted Proximal Supervised Fine-Tuning (TPSFT) loss over the editing dataset , while the rest of the model parameters remain frozen:
Here, is a hyperparameter that defines the clipping radius, which controls the size of the trust region. The probability ratio is the core of this objective and is defined as:
where is the probability of generating the ground-truth answer given the context from the model with updated FFN parameters, and is the corresponding probability from the reference policy. At the start of the editing process, this reference policy is the initial vanilla LLM. For each subsequent edit instance in the sequential editing process, it is then updated to be the state of the model resulting from the immediately preceding edit.
This objective function creates a trust-region constraint that is critical for balanced knowledge editing. The term aims to increase the likelihood of the correct answer, which is analogous to the objective in standard supervised fine-tuning. However, the ‘clip‘ function prevents this ratio from deviating too far from 1. When the updated model becomes significantly more confident about the target answer than the original model (i.e., when ), the gradient signal is effectively nullified for that instance. This mechanism acts as a powerful regularization, discouraging overly aggressive updates that could lead to overfitting on the new fact and, consequently, the disruption of pre-trained capabilities.
By integrating targeted parameter updates with a constrained optimization objective, TPSFT directly addresses the editing trilemma. Locality is achieved by physically confining the updates to the FFN layers, which are hypothesized to be the primary repositories of factual knowledge. Reliability is enforced by the supervised objective that maximizes the probability of the new fact. Finally, stability is preserved by the PSFT clipping mechanism, which prevents drastic policy shifts and ensures that the model’s behavior remains stable and consistent across a wide range of inputs beyond the specific edit.
4.2 Knowledge Consolidation via Group Relative Policy Optimization
After the TPSFT stage, the edited model has incorporated new facts at the parametric level. However, these parametric changes do not automatically propagate to the model’s reasoning capabilities. To bridge this gap, we introduce a consolidation step using Group Relative Policy Optimization (GRPO) that aligns the model’s inference-time behavior with the injected knowledge.
We formulate the consolidation as a reinforcement learning problem. Given a reasoning dataset containing queries that require reasoning over the edited facts, we optimize the edited model to generate trajectories that demonstrate both factual accuracy and reasoning consistency. The objective maximizes expected reward while constraining deviation from the post-TPSFT model:
where serves as the reference policy (the model after TPSFT), and controls the strength of regularization.
We optimize this objective using the GRPO algorithm with the following surrogate loss:
where is the importance ratio, and is the group-relative advantage computed from a batch of sampled trajectories.
The reward function evaluates multiple aspects of the generated trajectory:
where measures factual accuracy (whether the final answer matches ), enforces task-specific output format requirements, encourages concise outputs without extraneous tokens, and rewards internal reasoning coherence and alignment between intermediate steps and the final answer.
This consolidation step effectively integrates the parametric knowledge acquired through TPSFT into the model’s CoT-based inference policy, ensuring that the edited facts are not merely memorized but can be coherently utilized in complex reasoning tasks while maintaining locality on unrelated inputs.
5 Experiments
5.1 Experiment Settings
Datasets and Models. This work utilizes 1000 samples from each of three benchmark datasets, ZsRE Levy et al. (2017), COUNTERFACT Meng et al. (2022a), and QAEdit Yang et al. (2025b), to comprehensively evaluate the performance on knowledge editing tasks. We select two widely used LLMs, Llama-3-8B-Instruct Dubey et al. (2024) and Qwen-2.5-7B-Instruct Li et al. (2025a), as the base models for editing. For general ability evaluation, we use C-Eval Huang et al. (2023), CoQA Reddy et al. (2019), DROP Dua et al. (2019), SQuAD 2.0 Rajpurkar et al. (2018) and LogiQA Liu et al. (2020).
Baselines. We compare our method against two main categories: Parametric In-Place Editing methods (FT-M Zhang et al. (2024a), MEMIT Meng et al. (2022b), ALPHAEDIT Fang et al. (2024), MMKE Fu et al. (2025)) and External-Assisted Editing methods (WISE Wang et al. (2024a)). Parametric In-Place Editing methods are the main focus of this work, and we select the most representative methods in this category as baselines. For External-Assisted Editing methods, we select WISE as it is the SOTA method in this category.
Implementation Details. We conduct experiments using EasyEdit Xu et al. (2025) for evaluating various baselines, and employ the lm-evaluation-harness for assessing general capabilities. TPSFT is implemented through PSFT Zhu et al. (2025) for edit stage, while GRPO is built upon the EasyR1 Zheng et al. (2025b) for the consolidation stage. The specific hyperparameters are shown in Appendix A.1.
Evaluation Metrics. We evaluate our method along two principal axes: editing performance and general capability preservation. To assess editing performance, we employ the LLM-as-judge framework from Yang et al. (2025b); Gao et al. (2024); Gu et al. (2024b); Wang et al. (2024b), which mitigates the overestimation issue inherent in token-based metrics. In this framework, we leverage GPT-4.1 for a binary (correct/incorrect) evaluation of the model’s edited outputs to measure three key aspects: Reliability (edit success), Generalization (effectiveness on related inputs), and Locality (impact on unrelated inputs). To ensure that the editing process does not compromise the model’s broader abilities, we further evaluate its general capability preservation. To this end, we report Accuracy on the classification benchmarks C-Eval and LogiQA, alongside Exact Match (EM) and F1 scores for the question-answering datasets CoQA, DROP, and SQuAD 2.0. Details of real-world evaluation is in Appendix A.7
5.2 Main Results
Table 2 presents our evaluation of EtCon against existing baselines across three benchmarks under real-world lifelong editing evaluations. EtCon consistently outperforms all baselines across both model architectures. On Qwen-2.5-7B-Instruct, EtCon achieves 69.4% Reliability on ZsRE and 75.1% on QAEdit, surpassing the strongest baseline ALPHAEDIT by 53.5 and 75.1 percentage points respectively. Similar improvements occur on Llama-3-8B-Instruct, where EtCon reaches 73.5% Reliability on ZsRE versus FT-M’s 16.6%. Notably, EtCon maintains strong Generalization scores (60.8% on ZsRE, 63.0% on QAEdit for Qwen-2.5) while preserving acceptable Locality (24.2%-33.6%), confirming that our approach successfully preserves unrelated knowledge while performing targeted edits.
The local editing methods (MEMIT and ALPHAEDIT) fail catastrophically in lifelong editing. MEMIT collapses entirely on Qwen-2.5-7B-Instruct with near-zero performance across all metrics. ALPHAEDIT performs marginally better but remains highly unstable: it achieves 15.9% Reliability on ZsRE but completely fails on COUNTERFACT and QAEdit (0.0% across all metrics) for Qwen-2.5. Even when ALPHAEDIT reaches 61.0% Reliability on COUNTERFACT for Llama-3, its Locality drops to 16.1%, indicating severe knowledge disruption. This failure stems from destructive interference between sequential edits, where uncontrolled accumulation of weight deltas causes exponential growth in layer norms, leading to model collapse.
FT-M and WISE show improved stability over local editing methods but remain far below EtCon’s performance. FT-M achieves only 5.6% Reliability on ZsRE for Qwen-2.5 compared to EtCon’s 69.4%, while WISE performs even worse at 4.5%. On Llama-3, FT-M’s best result (27.9% on COUNTERFACT) still falls 39.2 percentage points below EtCon. These substantial performance gaps validate the effectiveness of our approach, which we attribute to two key design choices: the local editing in the edit stage preserves unrelated knowledge, and more crucially, the consolidation stage enables the reasoning network to effectively utilize the edited knowledge, thereby completing the critical final step in the knowledge editing pipeline. Extended experiments with up to 3,000 sequential edits demonstrating EtCon’s robustness in larger-scale lifelong editing scenarios are provided in Appendix A.3.
| ZsRE | COUNTERFACT | QAEdit | ||||||||||
| Method | Reli. | Gen. | Loc. | Reli. | Gen. | Loc. | Reli. | Gen. | Loc. | |||
Pre-edit | 4.4 | 3.2 | 28.5 | 1.0 | 0.5 | 36.9 | 9.8 | 10.1 | 36.2 | ||
| FT-M | 5.6 | 5.5 | 23.1 | 3.2 | 3.1 | 24.4 | 14.6 | 14.5 | 30.7 | |||
| MEMIT | 0.0 | 0.1 | 0.0 | 0.0 | 0.2 | 0.1 | 0.4 | 0.3 | 0.2 | |||
| ALPHAEDIT | 15.9 | 11.5 | 6.8 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | |||
| WISE | 4.5 | 3.3 | 19.1 | 1.4 | 1.5 | 31.0 | 7.1 | 9.7 | 16.9 | |||
| EtCon | 69.4 | 60.8 | 24.4 | 59.6 | 43.2 | 29.7 | 75.1 | 63.0 | 32.3 | |||
Pre-edit | 2.8 | 2.4 | 38.6 | 0.6 | 0.8 | 31.8 | 12.7 | 12.5 | 44.3 | ||
| FT-M | 16.6 | 15.5 | 29.3 | 27.9 | 18.6 | 10.5 | 34.1 | 33.2 | 30.1 | |||
| MEMIT | 0.1 | 0.1 | 0.0 | 0.3 | 0.7 | 0.4 | 0.2 | 0.7 | 0.0 | |||
| ALPHAEDIT | 18.7 | 14.0 | 6.3 | 61.0 | 43.8 | 16.1 | 18.2 | 14.9 | 7.5 | |||
| WISE | 4.3 | 3.1 | 2.2 | 1.3 | 0.8 | 31.3 | 8.1 | 13.3 | 0.9 | |||
| EtCon | 73.5 | 63.1 | 30.2 | 67.1 | 53.4 | 24.2 | 70.7 | 62.7 | 33.6 | |||
5.3 Analysis of Consolidation Stage
| DataSet | Metric | Base | FT-M |
MMKE |
ALPHA |
EtCon | ||||||
| Edited Knowledge | |||||||||||||||
| QAEdit | Reli. | 12.6 | 14.6 | 42.3 | 12.2 | 37.2 | 0.0 | 0.0 | 75.1 | ||||||
| Gen. | 13.9 | 14.5 | 34.1 | 10.4 | 31.4 | 0.0 | 0.0 | 63.0 | |||||||
| Loc. | 36.2 | 30.7 | 31.9 | 34.2 | 31.0 | 0.0 | 0.0 | 32.3 | |||||||
| General Capabilities | |||||||||||||||
| C-Eval | Acc. | 79.49 | 75.93 | 76.97 | 79.27 | 78.83 | 23.02 | 23.03 | 78.45 | ||||||
| CoQA | EM | 54.47 | 21.33 | 26.22 | 60.30 | 59.07 | 0.00 | 0.00 | 55.13 | ||||||
| F1 | 70.13 | 38.74 | 46.64 | 74.60 | 73.33 | 0.00 | 0.00 | 69.41 | |||||||
| DROP | EM | 2.21 | 2.37 | 2.79 | 10.30 | 8.46 | 0.00 | 0.00 | 2.52 | ||||||
| F1 | 9.94 | 13.31 | 14.59 | 24.32 | 21.59 | 0.00 | 0.00 | 8.60 | |||||||
| SQuAD | EM | 9.88 | 2.88 | 4.37 | 13.79 | 12.05 | 50.07 | 50.07 | 9.85 | ||||||
| F1 | 18.88 | 11.17 | 13.53 | 21.10 | 19.55 | 50.07 | 50.07 | 19.59 | |||||||
| LogiQA | Acc. | 38.71 | 37.02 | 37.79 | 41.01 | 39.17 | 21.81 | 21.81 | 38.40 | ||||||
To comprehensively evaluate the effectiveness of the proposed Consolidation phase, we conducted extensive experiments on the QAEdit dataset. We augmented three baseline knowledge editing methods (FT-M, MMKE, and ALPHAEDIT) with our Consolidation phase and compared their performance against our proposed EtCon method. As shown in Table 3, incorporating the Consolidation phase into FT-M and MMKE yields substantial improvements of 25-28% in Reliability and approximately 20% in Generalization metrics. These gains demonstrate that the Consolidation phase effectively bridges the gap between edited parametric knowledge and the model’s actual generation behavior, enabling successful knowledge utilization in real-world scenarios. Moreover, evaluations on multiple general-purpose benchmarks confirm that the Consolidation stage preserves the model’s general capabilities, with FT-M and MMKE maintaining their original performance levels and even exhibiting marginal improvements in certain cases. This preservation of general capabilities while enhancing editing performance validates the non-destructive nature of our consolidation mechanism.
However, the Consolidation phase cannot repair damage incurred during the editing stage. While FT-M with Consolidation achieves 5-8 percentage point improvements in EM and F1 scores on CoQA, these metrics remain substantially below the original model’s performance, highlighting the importance of careful knowledge updates during editing. MMKE’s design protects general capabilities but at the cost of reduced editing efficacy compared to EtCon. ALPHAEDIT exhibits model collapse after editing, which even the Consolidation phase cannot rectify. Overall, the Consolidation phase proves indispensable for knowledge editing, enabling effective generalization of newly edited knowledge while maintaining the model’s general capabilities. We further validate EtCon’s compatibility with reasoning-oriented architectures (DeepSeek-R1-Distill-Qwen-7B) in Appendix A.4. A detailed time efficiency analysis is provided in Appendix A.5.
5.4 ABLATION STUDIES
Stage Methods Reli. Gen. Loc. C-Eval CoQA sQuAD 2.0 Base - 0.6 0.8 31.8 50.82 78.20 29.52 Edit. w/ SFT 1.4 0.3 30.7 48.66 75.76 26.52 w/ TPSFT 3.3 1.8 30.2 50.07 78.52 34.60 Consolidate. w/o 56.1 22.4 24.7 - - - w/o 51.6 27.2 25.1 - - - Complete 67.1 53.4 24.2 - - -
We conduct a thorough ablation study on the COUNTERFACT dataset using Llama-3-8B-Instruct to isolate the individual contributions of each key component in our EtCon framework, with results presented in Table 4. In the Edit Stage, we compare our TPSFT against standard SFT. The results indicate that neither method alone is sufficient to enable reliable application of new knowledge, as reflected by low success and generalization scores. However, TPSFT demonstrates a clear advantage in preserving the model’s general capabilities, significantly mitigating the degradation observed with SFT. In the Consolidation Stage, building upon the TPSFT edit, we ablate components of our comprehensive reward function. Removing the cleanliness reward () causes a significant performance drop. Upon inspection, we find this encourages ”reward hacking,” where the model generates extraneous content to maximize its score, such as both old and new facts. The performance degrades more severely upon removing the consistency reward (), leading to catastrophic failures in reliability; for instance, the model might state the correct answer and then immediately contradict it. These findings confirm that our comprehensive reward function is critical for preventing such reward hacking and effectively steering the consolidation process toward reliable and coherent reasoning. For illustrative case studies, please refer to Appendix A.8.
6 Conclusion
In this paper, we identified that the critical gap between theoretical performance and practical effectiveness of knowledge editing methods stems from the absence of a consolidation stage that integrates parametric knowledge into the LLM’s actual generation behavior. To address this, we proposed the Edit-then-Consolidate (EtCon) framework, which combines Targeted Proximal Supervised Fine-Tuning (TPSFT) for precise knowledge editing with Group Relative Policy Optimization (GRPO) for effective consolidation. TPSFT updates targeted FFN weights within trust-region constraints to ensure reliable edits while preserving pretrained capabilities. Then GRPO aligns the edited knowledge with the model’s CoT-based inference policy through a comprehensive reward function. Our controlled experiments demonstrated the necessity of the knowledge consolidation stage, and comprehensive evaluations showed that EtCon significantly outperforms existing methods in lifelong editing scenarios, achieving efficient knowledge updates while maintaining locality and preserving model capabilities. These results suggest that explicitly decoupling editing and consolidation represents a promising paradigm for practical knowledge editing.
Appendix A Appendix
A.1 Additional Implementation Details
For our baseline experiments, we utilize the EasyEdit framework. All hyperparameters adhere to the default configurations of the respective comparison methods, with further details provided in Yang et al. (2025b); Qi et al. (2025). For our proposed EtCon method, we update only the FFN down-projection layers (mlp.down_proj) in layers 7–11 of Llama-3-8B-Instruct and layers 5–9 of Qwen2.5-7B-Instruct, following prior works Zhang et al. (2024a); Meng et al. (2022a); Geva et al. (2021); Qi et al. (2024) on knowledge editing. In the editing stage, we use AdamW with learning rate and set of TPSFT, using an edit batch size of 1 to perform 1,000 sequential single-sample edits on the same model instance. TPSFT is trained for 5 epochs with at most 6 update steps per edit using early stopping. In the consolidation stage, we optimize the inference-time policy with Group Relative Policy Optimization (GRPO). The comprehensive reward function in Equation (5) uses the following weight coefficients: for , for , for , and for . These weights were determined through extensive empirical experiments to balance factual accuracy with output quality. All specific hyperparameters are available in Table 5.
| Configuration | Value |
| Model Configuration | |
| Precision | BFloat16 |
| Max Prompt Length | 2k |
| Max Response Length | 2k |
| Training Hyperparameters | |
| Learning Rate | |
| Optimizer | AdamW (BF16 variant) |
| Global Batch Size | 64 |
| Rollout Batch Size | 256 |
| Micro Batch Size (Update) | 4 |
| Micro Batch Size (Experience) | 16 |
| Training Step | 100 |
| Gradient Clipping | 1.0 |
| Rollout Configuration | |
| Number of Rollouts () | 8 |
| Temperature | 1.0 |
| Top-p | 0.99 |
| Infrastructure | |
| GPUs | 8 × NVIDIA H800 |
| Tensor Parallelism | 1 |
| FSDP | Enabled |
| CPU Offloading | Disabled |
| Gradient Checkpointing | Enabled |
| Validation | |
| Validation Batch Size | 512 |
| Validation Frequency | Every 5 episodes |
| Validation before Training | Yes |
A.2 Effect of Edited Layers on EtCon
| Layers | Reli. | Gener. | Local. | |
Llama-3-8b-Instruct |
7-8-9-10-11 | 73.5 | 63.1 | 30.2 |
| 12-13-14-15-16 | 78.1 | 58.3 | 24.1 | |
| 17-18-19-20-21 | 76.7 | 53.2 | 17.3 |
We further investigate the impact of editing different FFN layers on EtCon’s performance through an ablation study, as shown in Table 6 and Figure 6. We observe that under identical hyperparameter settings, editing early layers (Layers 7-11) outperforms deeper layers (Layers 17-21) in both locality and generalization. Analyzing the ”high reward, low performance” phenomenon in deeper layers shown in Figure 6, we find that editing deeper layers is more prone to triggering Reward Hacking. Combining existing mechanistic interpretability research—which suggests that shallow layers mainly store factual knowledge while deeper layers handle information integration and reasoning—we posit that this stems from a misalignment between retained knowledge and updated information. Editing only deeper layers may cause retained knowledge in shallow layers to conflict with the injected knowledge in deeper layers. Faced with this cognitive conflict, the LLM likely adopts a speculative strategy to maximize rewards. This reward hacking leads to internal model confusion, thereby causing a decline in performance.
A.3 Lifelong Editing
To evaluate the robustness of our framework in larger-scale lifelong editing scenarios, we extend the ZsRE sequential editing experiments to 3,000 single-sample edits on distinct ZsRE examples. At each checkpoint (600, 1200, 1800, 2400, and 3000 edits), we first run the knowledge consolidation phase and then evaluate the edited LLM before proceeding with further edits, as illustrated in Figure 7. Across all three metrics, EtCon exhibits graceful degradation: starting from the first 600 edits, its Reliability and Generalization remain high and decrease only moderately as the stream length triples, while Locality fluctuates within a narrow band without signs of collapse. In contrast, FT-M begins with much lower scores and deteriorates rapidly as more edits accumulate, with Reliability and Generalization approaching near-zero and Locality dropping sharply. These results indicate that EtCon can sustain thousands of sequential edits and still substantially outperform the baseline under lifelong editing scenarios.
A.4 Evaluation on Reasoning-Oriented Architectures
| Layers | Reli. | Gener. | Local. | |
Deepseek-R1-distill-Qwen-7B
5-6-7-8-9 | 88.6 | 53.5 | 17.0 |
| 13-14-15-16-17 | 79.6 | 42.6 | 16.1 | |
| 23-24-25-26-27 | 83.0 | 52.9 | 7.3 |
To assess the robustness of our method on architectures with inherent reasoning capabilities, we extended our evaluation to DeepSeek-R1-Distill-Qwen-7B using 1,000 samples from the ZsRE. As reported in Table 7, our method maintains high efficacy on this reasoning-oriented architecture. As reported in Table 7, EtCon maintains high editing efficacy on this reasoning-oriented model. In particular, editing the shallow layers (Layers 5–9) yields the best trade-off, achieving 88.6% Reliability and 53.5% Generalization while preserving acceptable Locality (17.0%). This suggests that our Edit-then-Consolidate paradigm is compatible with the model’s intrinsic reasoning processes rather than disrupting them.
The green curve exhibits the fastest initial convergence (Steps 0-20). Since these layers are proximal to the output projection, the model can quickly maximize the reward by establishing a direct mapping to the target answer. However, similar to the phenomenon observed in Appendix A.2, this configuration plateaus at a lower reliability level. This suggests that modifying only the deep layers could lead the model to adopt a ”shortcut” strategy which may leave the model vulnerable to the cognitive conflict between retained shallow knowledge and updated deep injections. In contrast, the red curve displays a distinct ”warm-up” phase (Steps 0-15) followed by a sustained ascent to the highest performance plateau. This trajectory indicates that injecting knowledge into the shallow layers requires more optimization steps for the consolidate stage to propagate the changes through autoregressive generation pipeline. Crucially, this delay reflects a constructive consolidation process: the model is realigning its actual generation behavior with the updated parametric knowledge base. This results in convergence stability and reliability, demonstrating that our Edit-then-Consolidate paradigm is compatible with the intrinsic reasoning mechanisms of complex architectures.
A.5 Time Efficiency Analysis
| Methods | TPSFT (Ours) | AlphaEdit | MEMIT | GRACE | WISE | FT-M |
| Avg. Time / Edit | 6.01s | 7.39s | 7.78s | 3.02s | 2.68s | 0.61s |
In our earlier analysis, we argued that reliable knowledge editing in realistic settings requires both a knowledge editing stage and a knowledge-consolidation stage, and that any speed advantage should be predicated on reliability. Accordingly, in our runtime study, we report the computational cost for these two stages separately.
Editing stage. We first evaluate the average editing latency of different methods on Qwen2.5-7B-Instruct using the QAE dataset. For each method, we perform a sequence of 100 single-instance edits and measure the wall-clock time required to complete each individual edit. All methods are run under the same configuration. The results are summarized in Table 8. TPSFT attains an average editing time of 6.01 seconds, which is comparable to that of AlphaEdit (7.39 s) and MEMIT (7.78 s). GRACE (3.02 s) and WISE (2.68 s) exhibit lower latency, while FT-M yields the lowest runtime (0.61 s). Overall, TPSFT incurs a per-edit cost that is on par with representative parameter-editing baselines, indicating that its computational overhead at this stage remains moderate and compatible with practical deployment.
Consolidation stage. Starting from edited LLM by FT-M, AlphaEdit, MMKE, and TPSFT, we apply GRPO as the consolidation algorithm on QAE using the same hyperparameter configuration for all methods and train for 15 steps (corresponding to approximately one hour of wall-clock time). Figure 9 plots the comprehensive reward curves. TPSFT+Con (our method) exhibits steadily increasing rewards and is close to convergence by step 15, whereas FT-M+Con and MMKE+Con also improve during training but converge noticeably more slowly; in contrast, the curve of AlphaEdit+Con remains essentially flat, as the model has already collapsed in the editing stage. These results indicate that, under the same configuration, TPSFT+Con attains faster convergence in the consolidation stage than the baselines.
A.6 CoT Generation and Processing
To generate Chain-of-Thought reasoning paths during TPSFT training, we employ the following prompt template that encourages natural reasoning while ensuring convergence to target answers:
Design Rationale: This prompt template serves three critical functions in our TPSFT implementation:
1.
Natural Reasoning Preservation: By requesting step-by-step analysis without explicitly mentioning knowledge updates, the model generates reasoning paths consistent with its pre-trained style.
2.
Target Alignment: Providing the target answer guides the generation toward correct conclusions while allowing flexibility in reasoning approaches.
3.
Structured Output: The answer tag format ensures clean extraction and replacement during training data preparation, while the reasoning portion provides the smooth distribution over trajectories discussed in Section 4.1.
After generation, we enclose the reasoning path within <think></think> tags to explicitly demarcate the thought process, and overwrite the content inside the <answer></answer> tags with the verified new target fact, yielding training labels that combine natural reasoning patterns with an exact gold answer. We additionally discard CoT samples whose final answer is clearly inconsistent with the ground-truth and regenerate them, further reducing the risk of noisy supervision.
A.7 Real-world evaluation Details
In this work, we follow the design Yang et al. (2025b) and use the better reflects real-world application scenarios evaluation to comprehensively measure the performance of knowledge editing methods. Specifically, our evaluation process consists of three key stages:
(1) For Input: To assess the model’s ability to deeply integrate and apply new knowledge, our inputs include both factual questions and instructions that require multi-step reasoning. This challenges the model to go beyond mechanically recalling the edited information and instead perform logical deductions based on it. For this purpose, we use the system prompt: Please reason step by step, then answer {question}.
(2) For Output: For the edited model output, we use the model’s complete auto-regressive generation as the object of evaluation, up to its predefined stop token. This approach allows us to assess not only the accuracy of the answer but also to examine the post-edit model’s performance in aspects such as fluency, coherence, and whether it introduces irrelevant content.
(3) LLM-as-a-Judge Evaluation: To achieve a scalable and objective evaluation, we introduce a more powerful Large Language Model (LLM) to act as a ”judge.” This judge model makes its decision by comprehensively considering the original question, the ground-truth answer (Target), and the full answer content from the edited model, ultimately providing a binary (correct/incorrect) judgment. The full judge prompts are provided in Fig 4 and Fig. 5
A.8 Analysis of Reward Hacking Cases
Analysis of Reward Hacking Patterns: The two cases in Figures. 10 and 11 reveal distinct failure modes in the absence of proper reward design. In Figure 10, the model exhibits ”self-correction” behavior—correctly reasoning through the problem but then artificially inserting the target answer followed by an immediate correction. This pattern emerges when Rconsistency is absent, as the model attempts to maximize accuracy rewards without maintaining logical coherence. Figure 11 demonstrates ”answer hedging” where the model provides multiple answers to maximize the probability of including the correct one. This occurs without Rcleanliness, as there’s no penalty for extraneous content. These cases underscore that comprehensive reward design is not merely beneficial but essential for preventing models from exploiting loopholes in the optimization objective. The 11.0% and 15.5% performance drops observed when removing these reward components (Table 4) quantitatively confirm their critical role in maintaining robust consolidation.

### v2

@@
-EtCon: Edit-then-Consolidate for Reliable Knowledge Editing
+EtCon: Edit then Consolidate for Reliable Knowledge Editing
 Abstract
 Knowledge editing aims to update specific facts in large language models (LLMs) without full retraining.
-Prior efforts sought to tune the knowledge layers of LLMs, proving effective for making selective edits.
-However, a significant gap exists between their performance in controlled, teacher-forcing evaluations and their real-world effectiveness in lifelong learning scenarios, which greatly limits their practical applicability.
-This work’s empirical analysis reveals two recurring issues associated with this gap: (1) Most traditional methods lead the edited model to overfit to the new fact, thereby degrading pre-trained capabilities; (2) There is a critical absence of a knowledge consolidation stage, leaving new facts insufficiently integrated into LLMs’ inference-time behavior under autoregressive generation, thereby leading to a mismatch between parametric knowledge and actual generation behavior.
-To this end, we propose Edit-then-Consolidate, a novel knowledge editing paradigm that aims to bridge the gap between theoretical knowledge editing methods and their real-world applicability.
-Specifically, (1) our framework mitigates overfitting via Targeted Proximal Supervised Fine-Tuning (TPSFT) that localizes the edit via a trust-region objective to limit policy drift; (2) Then, a consolidation stage using Group Relative Policy Optimization (GRPO) aligns the edited knowledge with CoT-based inference policy by optimizing trajectory-level behavior under comprehensive reward signals.
-Extensive experiments demonstrate our framework consistently improves editing reliability and generalization under real-world evaluations, while better preserving locality and pre-trained capabilities.
+Prior efforts sought to tune the knowledge layers of LLMs, achieving improved performance in controlled, teacher-forced evaluations.
+However, they still encounter challenges in real-world autoregressive generation scenarios, which greatly limit their practical applicability.
+Our empirical analysis reveals two issues: (1) Most methods degrade pre-trained capabilities after injecting new knowledge; (2) They may exhibit a discrepancy between stored parametric knowledge and inference-time autoregressive generation behavior.
+To this end, we propose EtCon, an edit-then-consolidate paradigm that couples targeted edits with post-edit consolidation.
+Specifically, our framework comprises two stages: (1) Targeted Proximal Supervised Fine-Tuning (TPSFT) performs a constrained targeted edit to update parametric knowledge while controlling policy drift.
+(2) Group Relative Policy Optimization (GRPO) consolidates the edit by aligning autoregressive trajectories with the intended fact.
+Extensive experiments demonstrate that our EtCon improves editing reliability and real-world generalization, while better preserving pre-trained capabilities.
+[…]
 1 Introduction
-Large language models (LLMs) have demonstrated unprecedented capabilities across numerous tasks Guo et al.
-(2025); Wang et al.
-(2025b), serving as foundational reasoning engines for information retrieval Yang et al.
-(2025a), task automation agents He et al.
-(2025); Liu et al.
-(2025b); Wang et al.
-(2025a), and scientific research Rosen et al.
-(2025); Shmatko et al.
-(2025).
-However, as the external world continuously evolves, the static nature of LLMs’ pre-trained knowledge renders specific versions rapidly obsolete Zheng et al.
-(2025a).
-While retraining a large-scale LLM with updated knowledge could address this limitation, it requires substantial computational resources and pre-training data to maintain both knowledge update efficacy and general capabilities, making frequent knowledge updates impractical Mitchell et al.
-(2022).
-Knowledge editing methods Zhang et al.
-(2025); Scialanga et al.
-(2025); Li et al.
-(2025b); Rozner et al.
-(2024) have thus garnered significant attention as techniques that achieve targeted knowledge updates through localized parameter modifications while avoiding extensive resource consumption.
-Knowledge editing methods can be categorized into three main paradigms: (1) Parametric in-place editing methods Meng et al.
-(2022a); Dai et al.
-(2025); Han et al.
-(2024); Fu et al.
-(2025), which directly compute weight updates and apply them to the LLM’s weight matrices, encompassing approaches such as locate-then-edit, parameter-efficient fine-tuning, and model merging—all of which preserve model architecture without requiring additional modules; (2) Meta-learning-based methods Hartvigsen et al.
-(2023); Li et al.
-(2025b); Tan et al.
-(2023) that train auxiliary hypernetworks to predict weight updates for specific parameters to achieve knowledge editing objectives; (3) Memory-based methods Wang et al.
-(2024d; a) that store new knowledge in external modules and train LLMs to activate these modules during inference involving updated knowledge.
-While these methods demonstrate promise in constrained evaluation scenarios such as single editing and teacher-forcing evaluations, a significant performance gap emerges in more realistic auto-regressive evaluation and lifelong editing Jiang et al.
-(2024); Tan et al.
-(2023); Chen et al.
-(2024), leading recent research Gu et al.
-(2024a); Huang et al.
-(2024) to question the reliability and practical utility of existing knowledge editing methods.
-In this work, we conduct a comprehensive investigation into the root causes of this performance gap, focusing on Parametric In-Place Editing methods due to their high potential for practical application in lifelong learning scenarios.
-Through comprehensive empirical analysis, we identify two critical issues at the root of this gap.
-First, most of traditional methods cause edited models to overfit to newly introduced facts.
-This overfitting leads to excessive specialization of model parameters to editing examples, thereby degrading pre-trained general capabilities including robust reasoning, linguistic fluency, and robustness.
-Second, and more critically, a fundamental absence of a dedicated knowledge consolidation phase is observed.
-This omission results in new information being superficially encoded at the parametric level, failing to establish deep integration with the LLM’s inference-time behavior under autoregressive generation.
-This discordance manifests as a critical decoupling between knowledge representation and inferential activation.
-While the model successfully incorporates updated knowledge parametrically, it consistently fails to retrieve, activate, or apply this knowledge during autoregressive generation.
-Figure 1 illustrates the core challenge we address.
-When existing methods edit a fact (e.g., updating Michael Jordan’s nationality), the model may parametrically encode the new information but fail to consistently apply it during reasoning.
-This manifests as contradictory outputs where the model simultaneously acknowledges both old and new facts, revealing a fundamental misalignment between parametric knowledge and actual generation behavior.
-Our Edit-then-Consolidate framework resolves this by introducing a crucial consolidation stage that aligns the edited knowledge with the model’s inference-time policy.
-To address the limitation illustrated above, we propose Edit-then-Consolidate (EtCon), a two-stage knowledge-editing paradigm.
-In the first stage, we employ Targeted Proximal Supervised Fine-Tuning (TPSFT)—a refined variant of PSFT Zhu et al.
-(2025) that selectively updates only the FFN layers identified as knowledge repositories.
-This targeted approach, combined with trust-region constraints, ensures localized edits that preserve the model’s broader capabilities.
-In the second stage, we introduce a critical consolidation phase using Group Relative Policy Optimization (GRPO) to align the parametric knowledge with the model’s CoT-based inference policy through trajectory-level optimization under comprehensive reward signals.
+Large language models (LLMs) have demonstrated unprecedented capabilities across numerous tasks (Wang et al., 2025; Yang et al., 2025a; He et al., 2025; Liu et al., 2025b; Rosen et al., 2025; Shmatko et al., 2025).
+However, as the world continuously evolves, static pre-trained knowledge quickly becomes outdated (Zheng et al., 2025a).
+While continual pre-training can refresh knowledge, it is computationally expensive and impractical to perform frequently at scale (Mitchell et al., 2022), motivating knowledge editing methods that perform targeted updates through localized parameter modifications (Zhang et al., 2025; Scialanga et al., 2025; Li et al., 2025; Rozner et al., 2024).
+Existing knowledge editing methods roughly fall into two families: In-Place Parametric Edits (Meng et al., 2022a; Dai et al., 2025; Han et al., 2024; Fu et al., 2025; Jiang et al., 2025) that directly update model weights under the vanilla architecture, and External-Assisted Edits (Hartvigsen et al., 2023; Li et al., 2025; Tan et al., 2023; Wang et al., 2024d, a) that introduce auxiliary modules (e.g., hypernetworks or memory) to support updates at inference time.
+Despite encouraging results under single-edit and teacher-forced evaluations, their reliability in real-world autoregressive generation remains limited (Jiang et al., 2024; Tan et al., 2023; Chen et al., 2024; Gu et al., 2024a; Huang et al., 2024).
+In this work, our empirical analysis reveals two recurring issues.
+(1) Pre-trained capability degradation.
+Many existing methods perform edits by maximizing the likelihood of the new target on a small set of editing samples.
+Without explicit constraints or regularization, this few-shot objective can induce over-specialization to the editing examples, degrading pre-trained capabilities.
+In lifelong editing, as illustrated in Figure 1 (left), such drift may accumulate across lifelong edits and may even lead to model collapse (Gupta et al., 2024, 2025b).
+(2) Parametric knowledge–behavior mismatch in autoregressive generation.
+More critically, Figure 1 (right) shows that even when the edited fact is stored parametrically, it may not be consistently expressed along autoregressive trajectories.
+A key reason is the train-test mismatch between teacher forcing and autoregressive inference: edits are optimized on ground-truth prefixes, whereas real-world generation conditions on self-generated prefixes, inducing a prefix distribution shift that makes edited behavior fragile and prone to reversion.
+To address these limitations, we propose EtCon, a new edit-then-consolidate paradigm (Fig.
+4) that couples constrained targeted edits with post-edit consolidation to improve both editing reliability and pre-trained capability preservation.
+Our training involves a two-stage pipeline: Stage 1 (Edit).
+Targeted Proximal Supervised Fine-Tuning (TPSFT), a refined variant of PSFT (Zhu et al., 2025), performs a constrained targeted edit by updating only knowledge-relevant FFN layers, with trust-region constraints and self-generated CoT-based targets.
+This design localizes parameter updates and mitigates over-specialization, thereby preserving pre-trained capabilities.
+Stage 2 (Consolidate).
+GRPO consolidates the edit via trajectory-level optimization under reward signals.
+This stage aligns the parametric update with inference-time behavior, improving the robustness of the edited knowledge along autoregressive trajectories.
 We conduct extensive experiments on three datasets with Llama-3-8B-Instruct and Qwen2.5-7B-Instruct.
-Under auto-regressive generation with natural stopping and an LLM-as-a-judge protocol using GPT-4.1 (OpenAI (2025)), Edit-then-Consolidate improves editing reliability and generalization by 35%-50% over strong baselines.
+Under auto-regressive generation with natural stopping and an LLM-as-a-judge protocol (OpenAI, 2025), EtCon improves editing reliability and generalization by 35%-50% over strong baselines.
 It also significantly enhances locality while preserving critical pre-trained capabilities.
-Our contributions can be summarized as follows: (1) We identify that the absence of a knowledge-consolidation stage creates a critical knowledge-behavior misalignment, serving as the key bottleneck to the real-world applicability of knowledge-editing methods.
-(2) We propose Edit-then-Consolidate (EtCon): TPSFT for localized parametric edits, followed by GRPO for trajectory-level consolidation that aligns parametric knowledge with actual generation behavior.
-(3) Extensive experiments demonstrate that EtCon improves editing reliability and generalization by 40%–50%, strengthens locality, and preserves pre-trained capabilities under realistic evaluation settings.
+Our contributions are as follows:
+(1) We empirically show that incorporating a post-edit consolidation stage improves knowledge editing reliability under realistic scenarios.
+(2) We propose EtCon, a novel knowledge editing paradigm: TPSFT for localized parametric edits, followed by GRPO for trajectory-level consolidation that aligns parametric knowledge with actual generation behavior.
+(3) Extensive experiments demonstrate that EtCon improves editing reliability and generalization, while preserving pre-trained capabilities under realistic evaluation settings.
+| Method | Reli. | Gener. | Local. | |
+| Llama-3-8B-Instruct | Pre-Edit | 2.8 | 2.4 | 38.6 |
+| Pre-Edit(+GRPO) | 5.2 (+2.4) | 4.7 (+2.3) | 38.4 (-0.2) | |
+| FT-M | 16.6 | 15.5 | 29.3 | |
+| FT-M(+GRPO) | 62.9 (+46.3) | 52.7 (+37.2) | 24.9 (-4.4) | |
+| ALPHAEDIT | 18.7 | 14.0 | 6.3 | |
+| ALPHAEDIT(+GRPO) | 50.4 (+31.7) | 38.7 (+24.7) | 5.4 (-0.9) |
 2 Related Work
@@
 From the perspective of model architecture, the three paradigms introduced above can also be coarsely grouped into two families.
-Parametric in-place editing methods preserve the vanilla LLM architecture.
-The locate-then-edit paradigm Meng et al.
-(2022a); Dai et al.
-(2025); Li et al.
-(2024); Zhong et al.
-(2025); Zhang et al.
-(2024c) identifies knowledge locations within LLMs and modifies targeted parameters through gradient-based or analytical solutions.
-PEFT methods Zhu et al.
-(2020); Han et al.
-(2024); Wang et al.
-(2024c); Gupta et al.
-(2025) directly update model parameters via regularized gradient descent to achieve knowledge updates while constraining side effects Liu et al.
-(2025a).
+In-Place Parametric Edits methods preserve the vanilla LLM architecture.
+The locate-then-edit paradigm (Meng et al., 2022a; Dai et al., 2025; Li et al., 2024; Zhong et al., 2025; Zhang et al., 2024c) identifies knowledge locations within LLMs and modifies targeted parameters through gradient-based or analytical solutions.
+PEFT methods (Zhu et al., 2020; Han et al., 2024; Wang et al., 2024c; Gupta et al., 2025a) directly update model parameters via regularized gradient descent to achieve knowledge updates while constraining side effects (Liu et al., 2025a).
 These approaches seamlessly integrate with existing deployment infrastructure without additional inference latency.
-External-assisted editing methods rely on auxiliary modules for knowledge modification.
-Meta-learning approaches Tan et al.
-(2023); Hartvigsen et al.
-(2023); Li et al.
-(2025b) train hypernetworks to generate parameter updates, while memory-based methods Hartvigsen et al.
-(2023); Zhang et al.
-(2024b); Chen et al.
-(2024) encode knowledge in external modules that the LLM retrieves during inference.
+External-Assisted Edits methods rely on auxiliary modules for knowledge modification.
+Meta-learning approaches (Tan et al., 2023; Hartvigsen et al., 2023; Li et al., 2025) train hypernetworks to generate parameter updates, while memory-based methods (Hartvigsen et al., 2023; Zhang et al., 2024b; Chen et al., 2024) encode knowledge in external modules that the LLM retrieves during inference.
 Despite their superior performance in balancing reliability and locality, external methods introduce deployment complexity.
-Given these trade-offs, our work advances parametric in-place editing for lifelong knowledge editing scenarios.
+Given these trade-offs, our work advances In-Place Parametric Edits with a focus on more realistic lifelong knowledge editing scenarios.
 2.2 Evaluation of Knowledge editing methods
-Existing research Fang et al.
-(2024); Qi et al.
-(2025); Scialanga et al.
-(2025) predominantly evaluates the effectiveness of knowledge editing methods using a standard set of metrics.
-Reliability assesses the success rate of editing by calculating the percentage where .
-Generalization evaluates the model’s ability to generalize to new knowledge post-editing, measured by the percentage where when presented with rephrased queries pertaining to the new knowledge.
-Locality measures the extent to which editing a specific fact preserves the model’s responses to questions related to neighboring, unedited facts.
-Conventionally, the evaluation input typically consists of simple queries with identical prompt formats, without additional contextual information.
-For the output, edited models’ responses are often truncated to a target answer length or constrained by examples to match a specific target answer format.
-In generation, teacher forcing is frequently employed, feeding ground truth tokens as input during the decoding process.
+Existing work (Fang et al., 2024; Qi et al., 2025; Scialanga et al., 2025) typically evaluates knowledge editing with three standard metrics.
+Reliability measures edit success as the fraction of cases where .
+Generalization applies the same criterion to rephrased queries about the edited knowledge.
+Locality assesses whether the edit preserves responses to related, unedited facts.
+In these protocols, inputs are often templated queries with minimal context, outputs are constrained to a fixed answer format (e.g., via truncation or in-context exemplars), and decoding commonly uses teacher forcing with ground-truth prefixes.
 Recent studies, however, have highlighted the fragility of such evaluation paradigms.
 Consequently, this paper adopts a realistic evaluation approach for knowledge editing methods.
-Details of the real-world evaluation framework are in appendix A.7
-3 The Missing Consolidation Stage in Knowledge Editing
-Recent studies have revealed a critical performance gap in knowledge editing methods: while achieving high success rates under controlled teacher-forcing evaluation, these methods exhibit catastrophic failures in realistic auto-regressive settings.
-This stark discrepancy undermines their practical utility and raises fundamental questions about the effectiveness of current approaches.
-Through systematic investigation, we identify that this failure stems not from the editing mechanism itself, but from a fundamental architectural omission—the absence of a knowledge consolidation stage.
-We hypothesize that successful knowledge updating requires a two-stage process: (1) an initial parametric edit that injects new information into LLMs’ weights, followed by (2) a consolidation phase that integrates this knowledge into the LLMs’ inference-time behavior under autoregressive generation.
-Without consolidation, edited knowledge remains superficially encoded at the parametric level, failing to propagate to the model’s actual generation behavior.
-| Method | Reli. | Gener. | Local. | |
-Llama-3-8b-Instruct |
-Pre-Edit | 2.8 | 2.4 | 38.6 |
-| Pre-Edit(+GRPO) | 5.2 | 4.7 | 38.4 | |
-| FT-M | 16.6 | 15.5 | 29.3 | |
-| FT-M(+GRPO) | 62.9 | 52.7 | 24.9 | |
-| ALPHAEDIT | 18.7 | 14.0 | 6.3 | |
-| ALPHAEDIT(+GRPO) | 50.4 | 38.7 | 5.4 |
-To validate this hypothesis, we conduct controlled experiments augmenting existing editing methods with a consolidation mechanism.
-Table 1 presents compelling evidence: introducing Group Relative Policy Optimization (GRPO) as a post-editing consolidation step dramatically transforms performance.
-For FT-M, reliability surges from 16.6% to 62.9% on Llama3-8b-Instruct, while ALPHAEDIT improves from 18.7% to 50.4%.
-Crucially, these gains extend to generalization metrics, indicating that consolidation enables genuine knowledge integration rather than superficial memorization.
-The reward trajectories in Fig.
-2 further illuminate the consolidation dynamics.
-The monotonic increase demonstrates stable knowledge integration, where the model progressively aligns its reasoning behavior with the edited knowledge.
-Notably, applying GRPO directly to unedited models yields minimal improvements (Pre-Edit: 2.8% → 5.2%), confirming that consolidation requires prior parametric editing as a foundation.
-These findings establish a critical insight: the limitations of current knowledge editing methods arise from treating editing as a single-stage process.
-The Edit-then-Consolidate paradigm we propose addresses this fundamental gap, recognizing that parametric updates and behavioral alignment are complementary but distinct requirements for successful knowledge editing.
-4 The Edit-then-Consolidate Framework
-Building on the observational evidence in the preceding section, we posit that the limitations of current LLM knowledge-editing methods arise primarily from the lack of a principled consolidation stage that integrates edited knowledge with the model’s reasoning behavior; moreover, repeated overfitting edits can erode general abilities.
-To address this, we introduce Edit-then-Consolidate paradigm: Stage I employs Targeted Proximal Supervised Fine-Tuning (TPSFT) to perform localized knowledge editing under trust-region–style constraints, thereby limiting spillover while preserving pre-trained abilities; Stage II applies Group Relative Policy Optimization (GRPO) with a task-appropriate comprehensive reward to consolidate at the trajectory level under real-world evaluation signals.
-The remainder of this section presents the design rationale and the interaction between these two stages.
+The full real-world evaluation framework is described in Appendix A.6.
+3 Preliminary
+3.1 Group Relative Policy Optimization
+Group Relative Policy Optimization (GRPO) (Guo et al., 2025) is a trajectory-level reinforcement learning algorithm that optimizes language model policies through relative reward comparisons within sampled groups.
+Given a policy and a dataset , GRPO maximizes expected reward while constraining deviation from a reference policy :
+where is a reward function and controls regularization strength. The optimization employs a clipped surrogate objective:
+where is the importance ratio, and is the group-relative advantage computed from trajectories sampled for the same query .
+This design eliminates the need for a separate critic network while enabling stable policy updates.
+Notably, GRPO’s trajectory-level optimization makes it naturally suited for aligning knowledge that is parametrically encoded but not reliably applied during autoregressive generation.
+3.2 Motivation for Knowledge Consolidation
+Our key insight is that successful knowledge editing requires a two-stage process rather than parametric knowledge editing alone.
+Specifically, recent studies reveal a critical gap: while existing methods achieve high success under teacher-forcing evaluation, they can fail in realistic autoregressive generation.
+Through systematic investigation, we posit that an explicit knowledge consolidation stage is crucial for mitigating this gap.
+Without a knowledge consolidation stage, edited knowledge may remains superficially encoded at the parametric level, failing to propagate to the model’s generation behavior.
+To validate this hypothesis, we augment existing editing methods with GRPO as a post-editing consolidation step.
+As shown in Table 3 and Figure 3, introducing consolidation dramatically improves both reliability and generalization across multiple baselines.
+Furthermore, applying GRPO directly to unedited models yields minimal gains, confirming that consolidation requires prior parametric editing as a foundation.
+These findings establish that parametric knowledge updates and behavioral alignment are complementary but distinct requirements for successful knowledge editing.
+4 EtCon
+In this section, we present the EtCon framework (as shown in Figure 4).
+Stage I employs Targeted Proximal Supervised Fine-Tuning (TPSFT) to perform localized knowledge editing with PPO-style ratio clipping that constrains policy divergence, thereby limiting spillover while preserving pre-trained abilities (§4.1); Stage II applies GRPO with a comprehensive reward to consolidate edited knowledge at the trajectory level under autoregressive generation (§4.2).
 4.1 Knowledge Editing via Targeted Proximal Fine-Tuning
-In this section, we introduce Targeted Proximal Supervised Fine-Tuning (TPSFT) as a refined knowledge-editing method that addresses the trilemma of reliability, locality, and generality.
-This approach differs from raw PSFT that update the whole LLMs, by selectively update only the FFNs of LLMs.
-This targeted update strategy effectively injects new knowledge while minimizing disruption to the model’s overall architecture and pre-trained capabilities.
-We consider a knowledge editing dataset that contains editing instances , where each context contains a question about the new fact, and is the corresponding ground-truth answer.
-At the start of the editing process, we have a vanilla LLM parameterized by .
-We partition the model’s parameters into two disjoint sets: the target FFN parameters to be edited, , and the remaining frozen parameters, , such that .
-The objective of TPSFT is to learn a new set of FFN parameters, , yielding an updated model where .
-This model must accurately generate the target answer for a given context , while minimizing disruptions to its performance on unrelated inputs.
-A critical innovation in our TPSFT is the use of CoT-augmented training labels.
-For each editing instance , we: (1) prompt the vanilla LLM to generate a CoT reasoning path for using an instruction template (see Appendix), and (2) replace the generated answer with the target new fact , yielding the training label .
-This design enables learning smoothed distributions over reasoning paths rather than sharp one-hot targets.
-More importantly, it preserves the model’s natural reasoning patterns—the model learns to reach new answers through its inherent reasoning style rather than abandoning pre-trained capabilities.
-This significantly reduces disruption while ensuring accurate knowledge updates.
-To achieve this, we update the targeted FFN parameters by minimizing the following Targeted Proximal Supervised Fine-Tuning (TPSFT) loss over the editing dataset , while the rest of the model parameters remain frozen:
-Here, is a hyperparameter that defines the clipping radius, which controls the size of the trust region. The probability ratio is the core of this objective and is defined as:
-where is the probability of generating the ground-truth answer given the context from the model with updated FFN parameters, and is the corresponding probability from the reference policy.
+In this section, we introduce Targeted Proximal Supervised Fine-Tuning (TPSFT) as a refined knowledge-editing method designed to improve editing efficacy while better preserving pre-trained capabilities.
+This approach differs from raw PSFT that updates the whole LLMs, we restrict updates exclusively to specific FFN layers, widely recognized as the primary stores of factual knowledge (Meng et al., 2022a).
+This targeted update strategy effectively injects new knowledge without modifying the model architecture, while minimizing disruption to pre-trained capabilities.
+We consider a knowledge editing instance , where is the input query and is the target answer.
+Let parameterized by denote the vanilla LLM before the current edit.
+We partition the current parameters into two disjoint sets: the trainable FFN parameters and the frozen parameters , such that .
+The objective of TPSFT is to optimize to learn the updated parameters , yielding the post-edit policy where .
+A distinctive aspect of TPSFT is the use of self-generated CoT-augmented training labels.
+For each instance , we: (1) prompt the vanilla LLM to generate a CoT reasoning path for (see Appendix A.5), and (2) replace the generated answer with , yielding the training target .
+This design integrates the CoT into the supervision signal.
+Rather than overfitting to a sharp, isolated input-output mapping, the model learns to associate the new knowledge with valid reasoning patterns.
+To ensure data quality, we filter out generated chains that are logically incompatible with the target (see Appendix A.5), thereby preserving the model’s intrinsic generation capabilities while injecting the new fact.
+Formally, given the editing dataset , for each token position in the target sequence , we define the state and the target token . We optimize by minimizing the following TPSFT loss:
+Here, is a hyperparameter that defines the clipping threshold.
+Following Zhu et al.
+(2025), the ratio clipping mechanism provides implicit trust-region regularization by (i) bounding the log-probability change per sample to , and (ii) nullifying gradients when the ratio exceeds the upper bound.
+The probability ratio is the core of this objective and is defined as:
+where is the probability of generating the target token given the context from the model with the current FFN parameters, and is the corresponding probability from the reference policy.
 At the start of the editing process, this reference policy is the initial vanilla LLM.
 For each subsequent edit instance in the sequential editing process, it is then updated to be the state of the model resulting from the immediately preceding edit.
-This objective function creates a trust-region constraint that is critical for balanced knowledge editing.
-The term aims to increase the likelihood of the correct answer, which is analogous to the objective in standard supervised fine-tuning.
-However, the ‘clip‘ function prevents this ratio from deviating too far from 1.
-When the updated model becomes significantly more confident about the target answer than the original model (i.e., when ), the gradient signal is effectively nullified for that instance.
-This mechanism acts as a powerful regularization, discouraging overly aggressive updates that could lead to overfitting on the new fact and, consequently, the disruption of pre-trained capabilities.
-By integrating targeted parameter updates with a constrained optimization objective, TPSFT directly addresses the editing trilemma.
-Locality is achieved by physically confining the updates to the FFN layers, which are hypothesized to be the primary repositories of factual knowledge.
-Reliability is enforced by the supervised objective that maximizes the probability of the new fact.
-Finally, stability is preserved by the PSFT clipping mechanism, which prevents drastic policy shifts and ensures that the model’s behavior remains stable and consistent across a wide range of inputs beyond the specific edit.
-4.2 Knowledge Consolidation via Group Relative Policy Optimization
-After the TPSFT stage, the edited model has incorporated new facts at the parametric level.
-However, these parametric changes do not automatically propagate to the model’s reasoning capabilities.
-To bridge this gap, we introduce a consolidation step using Group Relative Policy Optimization (GRPO) that aligns the model’s inference-time behavior with the injected knowledge.
-We formulate the consolidation as a reinforcement learning problem.
-Given a reasoning dataset containing queries that require reasoning over the edited facts, we optimize the edited model to generate trajectories that demonstrate both factual accuracy and reasoning consistency.
-The objective maximizes expected reward while constraining deviation from the post-TPSFT model:
-where serves as the reference policy (the model after TPSFT), and controls the strength of regularization.
-We optimize this objective using the GRPO algorithm with the following surrogate loss:
-where is the importance ratio, and is the group-relative advantage computed from a batch of sampled trajectories.
-The reward function evaluates multiple aspects of the generated trajectory:
+This objective function acts as a trust-region regularizer via ratio clipping.
+The term encourages the model to increase the likelihood of the target answer, similar to standard supervised fine-tuning.
+Crucially, when this ratio exceeds the threshold , the gradient with respect to is nullified.
+This design effectively removes optimization pressure from well-learned samples, constraining parametric deviation and preventing individual edits from dominating the parameter updates.
+However, despite these constraints, TPSFT remains a teacher-forced training mechanism implicitly conditioned on ground-truth prefixes .
+During inference, the model must rely on its own generated history .
+This discrepancy creates a fundamental train-inference mismatch: while the knowledge is parametrically encoded, it may not be reliably elicited during autoregressive generation.
+This critical limitation necessitates the subsequent consolidation phase.
+4.2 Knowledge Consolidation via GRPO
+As motivated in § 3.2, parametric editing alone is insufficient—edited knowledge must be consolidated at the behavioral level.
+Building on GRPO’s trajectory-level optimization (§ 3.1), we design a consolidation algorithm that transforms parametrically encoded knowledge into reliably generated outputs.
+Starting from the edited model obtained via TPSFT, we optimize the policy (initialized as ) through autoregressive sampling on the editing dataset .
+For each query , the model generates complete trajectories , which are then scored against the target answer .
+The group-relative advantage formulation enables the model to learn from comparative trajectory quality without requiring absolute reward calibration.
+The reference policy is set to , ensuring that consolidation refines generation behavior while preserving the parametric edits from Stage I.
+| ZsRE | COUNTERFACT | QAEdit | |||||||||||
+| Method | Reli. | Gen. | Loc. | Avg. | Reli. | Gen. | Loc. | Avg. | Reli. | Gen. | Loc. | Avg. | |
+| Qwen2.5-7B -Instruct | Pre-edit | 4.4 | 3.2 | 28.5 | 12.0 | 1.0 | 0.5 | 36.9 | 12.8 | 9.8 | 10.1 | 36.2 | 18.7 |
+| FT-M | 5.6 | 5.5 | 23.1 | 11.4 | 3.2 | 3.1 | 24.4 | 10.2 | 14.6 | 14.5 | 30.7 | 19.9 | |
+| MEMIT | 0.0 | 0.1 | 0.0 | 0.0 | 0.0 | 0.2 | 0.1 | 0.1 | 0.4 | 0.3 | 0.2 | 0.3 | |
+| ALPHAEDIT | 15.9 | 11.5 | 6.8 | 11.4 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | |
+| WISE | 4.5 | 3.3 | 19.1 | 9.0 | 1.4 | 1.5 | 31.0 | 11.3 | 7.1 | 9.7 | 16.9 | 11.2 | |
+| GRACE | 4.8 | 4.7 | 28.1 | 12.5 | 1.9 | 1.2 | 36.8 | 13.3 | 14.2 | 13.0 | 36.0 | 21.1 | |
+[…]
+| Llama-3 -8B-Instruct | Pre-edit | 2.8 | 2.4 | 38.6 | 14.6 | 0.6 | 0.8 | 31.8 | 11.1 | 12.7 | 12.5 | 44.3 | 23.2 |
+| FT-M | 16.6 | 15.5 | 29.3 | 20.5 | 27.9 | 18.6 | 10.5 | 19.0 | 34.1 | 33.2 | 30.1 | 32.5 | |
+| MEMIT | 0.1 | 0.1 | 0.0 | 0.1 | 0.3 | 0.7 | 0.4 | 0.5 | 0.2 | 0.7 | 0.0 | 0.3 | |
+| ALPHAEDIT | 18.7 | 14.0 | 6.3 | 13.0 | 61.0 | 43.8 | 16.1 | 40.3 | 18.2 | 14.9 | 7.5 | 13.5 | |
+| WISE | 4.3 | 3.1 | 2.2 | 3.2 | 1.3 | 0.8 | 31.3 | 11.1 | 8.1 | 13.3 | 0.9 | 7.4 | |
+| GRACE | 5.5 | 3.9 | 37.9 | 15.8 | 2.1 | 0.7 | 32.1 | 11.6 | 15.4 | 13.4 | 34.2 | 21.0 | |
+[…]
+The reward function evaluates multiple aspects of the generated trajectory (as illustrated in Figure 4):
 where measures factual accuracy (whether the final answer matches ), enforces task-specific output format requirements, encourages concise outputs without extraneous tokens, and rewards internal reasoning coherence and alignment between intermediate steps and the final answer.
-This consolidation step effectively integrates the parametric knowledge acquired through TPSFT into the model’s CoT-based inference policy, ensuring that the edited facts are not merely memorized but can be coherently utilized in complex reasoning tasks while maintaining locality on unrelated inputs.
+This consolidation step integrates the parametric knowledge acquired through TPSFT into the model’s inference policy.
+By optimizing trajectory-level generation under autoregressive sampling rather than teacher-forcing, GRPO mitigates the train-inference mismatch, ensuring that edited knowledge is reliably elicited in actual behavior.
 5 Experiments
@@
 Datasets and Models.
-This work utilizes 1000 samples from each of three benchmark datasets, ZsRE Levy et al.
-(2017), COUNTERFACT Meng et al.
-(2022a), and QAEdit Yang et al.
-(2025b), to comprehensively evaluate the performance on knowledge editing tasks.
-We select two widely used LLMs, Llama-3-8B-Instruct Dubey et al.
-(2024) and Qwen-2.5-7B-Instruct Li et al.
-(2025a), as the base models for editing.
-For general ability evaluation, we use C-Eval Huang et al.
-(2023), CoQA Reddy et al.
-(2019), DROP Dua et al.
-(2019), SQuAD 2.0 Rajpurkar et al.
-(2018) and LogiQA Liu et al.
-(2020).
+This work utilizes 1000 samples from each of four benchmark, ZsRE (Levy et al., 2017), COUNTERFACT (Meng et al., 2022a), MQuAKE-CF-v2 (Zhong et al., 2023) (1-edit subset), and QAEdit (Yang et al., 2025b), to evaluate the performance on knowledge editing tasks.
+We focus on the lifelong editing with sequential edits (batch size 1), which better reflects realistic deployment scenarios where knowledge updates arrive continuously over time.
+We select the Llama-3-8B-Instruct (Dubey et al., 2024) and Qwen2.5-7B-Instruct (Qwen et al., 2025), as the base models for editing.
+For general ability evaluation, we use C-Eval (Huang et al., 2023), CoQA (Reddy et al., 2019), DROP (Dua et al., 2019), SQuAD 2.0 (Rajpurkar et al., 2018) and LogiQA (Liu et al., 2020).
 Baselines.
-We compare our method against two main categories: Parametric In-Place Editing methods (FT-M Zhang et al.
-(2024a), MEMIT Meng et al.
-(2022b), ALPHAEDIT Fang et al.
-(2024), MMKE Fu et al.
-(2025)) and External-Assisted Editing methods (WISE Wang et al.
-(2024a)).
-Parametric In-Place Editing methods are the main focus of this work, and we select the most representative methods in this category as baselines.
-For External-Assisted Editing methods, we select WISE as it is the SOTA method in this category.
+We compare our method against two main categories: In-Place Parametric Edits methods (FT-M (Zhang et al., 2024a), MEMIT (Meng et al., 2022b), ALPHAEDIT (Fang et al., 2024), MMKE (Fu et al., 2025)) and External-Assisted Edits methods (WISE (Wang et al., 2024a), GRACE (Hartvigsen et al., 2023)).
+In-Place Parametric Edits methods are the main focus of this work, and we select the most representative methods as baselines.
 Implementation Details.
-We conduct experiments using EasyEdit Xu et al.
-(2025) for evaluating various baselines, and employ the lm-evaluation-harness for assessing general capabilities.
-TPSFT is implemented through PSFT Zhu et al.
-(2025) for edit stage, while GRPO is built upon the EasyR1 Zheng et al.
-(2025b) for the consolidation stage.
-The specific hyperparameters are shown in Appendix A.1.
+We conduct experiments using EasyEdit (Xu et al., 2025) for evaluating various baselines, and employ the lm-evaluation-harness (Gao et al., 2024a) for assessing general capabilities.
+TPSFT is implemented through PSFT (Zhu et al., 2025) for edit stage, while GRPO is built upon the EasyR1 (Zheng et al., 2025b) for the consolidation stage.
+More details are in Appendix A.1.
 Evaluation Metrics.
-We evaluate our method along two principal axes: editing performance and general capability preservation.
-To assess editing performance, we employ the LLM-as-judge framework from Yang et al.
-(2025b); Gao et al.
-(2024); Gu et al.
-(2024b); Wang et al.
-(2024b), which mitigates the overestimation issue inherent in token-based metrics.
-In this framework, we leverage GPT-4.1 for a binary (correct/incorrect) evaluation of the model’s edited outputs to measure three key aspects: Reliability (edit success), Generalization (effectiveness on related inputs), and Locality (impact on unrelated inputs).
-To ensure that the editing process does not compromise the model’s broader abilities, we further evaluate its general capability preservation.
-To this end, we report Accuracy on the classification benchmarks C-Eval and LogiQA, alongside Exact Match (EM) and F1 scores for the question-answering datasets CoQA, DROP, and SQuAD 2.0.
-Details of real-world evaluation is in Appendix A.7
+We evaluate our method along two principal axes: editing performance and pre-trained capability preservation.
+To assess editing performance, we employ the LLM-as-judge framework from (Yang et al., 2025b; Gao et al., 2024b; Gu et al., 2024b; Wang et al., 2024b), which mitigates the overestimation issue inherent in token-based metrics.
+Note that while GRPO training uses efficient rule-based rewards (§4.2), evaluation employs GPT-4.1 for semantic assessment to ensure rigor.
+Specifically, we conduct a binary (correct/incorrect) evaluation of the model’s edited outputs.
+For ZsRE, COUNTERFACT, and QAEdit, we measure Reliability (edit success), Generalization (effectiveness on related inputs), and Locality (impact on unrelated inputs).
+For MQuAKE-CF-v2, we report Edit-wise accuracy and Multi-hop accuracy.
+We report pretrained capability on C-Eval and LogiQA, and Exact Match (EM) and F1 on CoQA, DROP, and SQuAD 2.0.
+Real-world evaluation details are provided in Appendix A.6.
 5.2 Main Results
-Table 2 presents our evaluation of EtCon against existing baselines across three benchmarks under real-world lifelong editing evaluations.
-EtCon consistently outperforms all baselines across both model architectures.
-On Qwen-2.5-7B-Instruct, EtCon achieves 69.4% Reliability on ZsRE and 75.1% on QAEdit, surpassing the strongest baseline ALPHAEDIT by 53.5 and 75.1 percentage points respectively.
-Similar improvements occur on Llama-3-8B-Instruct, where EtCon reaches 73.5% Reliability on ZsRE versus FT-M’s 16.6%.
-Notably, EtCon maintains strong Generalization scores (60.8% on ZsRE, 63.0% on QAEdit for Qwen-2.5) while preserving acceptable Locality (24.2%-33.6%), confirming that our approach successfully preserves unrelated knowledge while performing targeted edits.
-The local editing methods (MEMIT and ALPHAEDIT) fail catastrophically in lifelong editing.
-MEMIT collapses entirely on Qwen-2.5-7B-Instruct with near-zero performance across all metrics.
-ALPHAEDIT performs marginally better but remains highly unstable: it achieves 15.9% Reliability on ZsRE but completely fails on COUNTERFACT and QAEdit (0.0% across all metrics) for Qwen-2.5.
-Even when ALPHAEDIT reaches 61.0% Reliability on COUNTERFACT for Llama-3, its Locality drops to 16.1%, indicating severe knowledge disruption.
-This failure stems from destructive interference between sequential edits, where uncontrolled accumulation of weight deltas causes exponential growth in layer norms, leading to model collapse.
-FT-M and WISE show improved stability over local editing methods but remain far below EtCon’s performance.
-FT-M achieves only 5.6% Reliability on ZsRE for Qwen-2.5 compared to EtCon’s 69.4%, while WISE performs even worse at 4.5%.
-On Llama-3, FT-M’s best result (27.9% on COUNTERFACT) still falls 39.2 percentage points below EtCon.
-These substantial performance gaps validate the effectiveness of our approach, which we attribute to two key design choices: the local editing in the edit stage preserves unrelated knowledge, and more crucially, the consolidation stage enables the reasoning network to effectively utilize the edited knowledge, thereby completing the critical final step in the knowledge editing pipeline.
-Extended experiments with up to 3,000 sequential edits demonstrating EtCon’s robustness in larger-scale lifelong editing scenarios are provided in Appendix A.3.
-| ZsRE | COUNTERFACT | QAEdit | ||||||||||
-| Method | Reli. | Gen. | Loc. | Reli. | Gen. | Loc. | Reli. | Gen. | Loc. | |||
-Pre-edit | 4.4 | 3.2 | 28.5 | 1.0 | 0.5 | 36.9 | 9.8 | 10.1 | 36.2 | ||
-| FT-M | 5.6 | 5.5 | 23.1 | 3.2 | 3.1 | 24.4 | 14.6 | 14.5 | 30.7 | |||
-| MEMIT | 0.0 | 0.1 | 0.0 | 0.0 | 0.2 | 0.1 | 0.4 | 0.3 | 0.2 | |||
-| ALPHAEDIT | 15.9 | 11.5 | 6.8 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | |||
-| WISE | 4.5 | 3.3 | 19.1 | 1.4 | 1.5 | 31.0 | 7.1 | 9.7 | 16.9 | |||
-| EtCon | 69.4 | 60.8 | 24.4 | 59.6 | 43.2 | 29.7 | 75.1 | 63.0 | 32.3 | |||
-Pre-edit | 2.8 | 2.4 | 38.6 | 0.6 | 0.8 | 31.8 | 12.7 | 12.5 | 44.3 | ||
-| FT-M | 16.6 | 15.5 | 29.3 | 27.9 | 18.6 | 10.5 | 34.1 | 33.2 | 30.1 | |||
-| MEMIT | 0.1 | 0.1 | 0.0 | 0.3 | 0.7 | 0.4 | 0.2 | 0.7 | 0.0 | |||
-| ALPHAEDIT | 18.7 | 14.0 | 6.3 | 61.0 | 43.8 | 16.1 | 18.2 | 14.9 | 7.5 | |||
-| WISE | 4.3 | 3.1 | 2.2 | 1.3 | 0.8 | 31.3 | 8.1 | 13.3 | 0.9 | |||
-| EtCon | 73.5 | 63.1 | 30.2 | 67.1 | 53.4 | 24.2 | 70.7 | 62.7 | 33.6 | |||
-5.3 Analysis of Consolidation Stage
-| DataSet | Metric | Base | FT-M |
-MMKE |
-ALPHA |
-EtCon | ||||||
-| Edited Knowledge | |||||||||||||||
-| QAEdit | Reli. | 12.6 | 14.6 | 42.3 | 12.2 | 37.2 | 0.0 | 0.0 | 75.1 | ||||||
-| Gen. | 13.9 | 14.5 | 34.1 | 10.4 | 31.4 | 0.0 | 0.0 | 63.0 | |||||||
-| Loc. | 36.2 | 30.7 | 31.9 | 34.2 | 31.0 | 0.0 | 0.0 | 32.3 | |||||||
-| General Capabilities | |||||||||||||||
-| C-Eval | Acc. | 79.49 | 75.93 | 76.97 | 79.27 | 78.83 | 23.02 | 23.03 | 78.45 | ||||||
-| CoQA | EM | 54.47 | 21.33 | 26.22 | 60.30 | 59.07 | 0.00 | 0.00 | 55.13 | ||||||
-| F1 | 70.13 | 38.74 | 46.64 | 74.60 | 73.33 | 0.00 | 0.00 | 69.41 | |||||||
-| DROP | EM | 2.21 | 2.37 | 2.79 | 10.30 | 8.46 | 0.00 | 0.00 | 2.52 | ||||||
-| F1 | 9.94 | 13.31 | 14.59 | 24.32 | 21.59 | 0.00 | 0.00 | 8.60 | |||||||
-| SQuAD | EM | 9.88 | 2.88 | 4.37 | 13.79 | 12.05 | 50.07 | 50.07 | 9.85 | ||||||
-| F1 | 18.88 | 11.17 | 13.53 | 21.10 | 19.55 | 50.07 | 50.07 | 19.59 | |||||||
-| LogiQA | Acc. | 38.71 | 37.02 | 37.79 | 41.01 | 39.17 | 21.81 | 21.81 | 38.40 | ||||||
-To comprehensively evaluate the effectiveness of the proposed Consolidation phase, we conducted extensive experiments on the QAEdit dataset.
-We augmented three baseline knowledge editing methods (FT-M, MMKE, and ALPHAEDIT) with our Consolidation phase and compared their performance against our proposed EtCon method.
-As shown in Table 3, incorporating the Consolidation phase into FT-M and MMKE yields substantial improvements of 25-28% in Reliability and approximately 20% in Generalization metrics.
-These gains demonstrate that the Consolidation phase effectively bridges the gap between edited parametric knowledge and the model’s actual generation behavior, enabling successful knowledge utilization in real-world scenarios.
-Moreover, evaluations on multiple general-purpose benchmarks confirm that the Consolidation stage preserves the model’s general capabilities, with FT-M and MMKE maintaining their original performance levels and even exhibiting marginal improvements in certain cases.
-This preservation of general capabilities while enhancing editing performance validates the non-destructive nature of our consolidation mechanism.
-However, the Consolidation phase cannot repair damage incurred during the editing stage.
-While FT-M with Consolidation achieves 5-8 percentage point improvements in EM and F1 scores on CoQA, these metrics remain substantially below the original model’s performance, highlighting the importance of careful knowledge updates during editing.
-MMKE’s design protects general capabilities but at the cost of reduced editing efficacy compared to EtCon.
-ALPHAEDIT exhibits model collapse after editing, which even the Consolidation phase cannot rectify.
-Overall, the Consolidation phase proves indispensable for knowledge editing, enabling effective generalization of newly edited knowledge while maintaining the model’s general capabilities.
-We further validate EtCon’s compatibility with reasoning-oriented architectures (DeepSeek-R1-Distill-Qwen-7B) in Appendix A.4.
-A detailed time efficiency analysis is provided in Appendix A.5.
-5.4 ABLATION STUDIES
-Stage Methods Reli.
-Gen.
-Loc.
-C-Eval CoQA sQuAD 2.0 Base - 0.6 0.8 31.8 50.82 78.20 29.52 Edit.
-w/ SFT 1.4 0.3 30.7 48.66 75.76 26.52 w/ TPSFT 3.3 1.8 30.2 50.07 78.52 34.60 Consolidate.
-w/o 56.1 22.4 24.7 - - - w/o 51.6 27.2 25.1 - - - Complete 67.1 53.4 24.2 - - -
-We conduct a thorough ablation study on the COUNTERFACT dataset using Llama-3-8B-Instruct to isolate the individual contributions of each key component in our EtCon framework, with results presented in Table 4.
-In the Edit Stage, we compare our TPSFT against standard SFT.
-The results indicate that neither method alone is sufficient to enable reliable application of new knowledge, as reflected by low success and generalization scores.
-However, TPSFT demonstrates a clear advantage in preserving the model’s general capabilities, significantly mitigating the degradation observed with SFT.
-In the Consolidation Stage, building upon the TPSFT edit, we ablate components of our comprehensive reward function.
-Removing the cleanliness reward () causes a significant performance drop.
-Upon inspection, we find this encourages ”reward hacking,” where the model generates extraneous content to maximize its score, such as both old and new facts.
-The performance degrades more severely upon removing the consistency reward (), leading to catastrophic failures in reliability; for instance, the model might state the correct answer and then immediately contradict it.
-These findings confirm that our comprehensive reward function is critical for preventing such reward hacking and effectively steering the consolidation process toward reliable and coherent reasoning.
-For illustrative case studies, please refer to Appendix A.8.
+Edit Efficiency.
+EtCon consistently achieves the best overall performance under real-world lifelong editing evaluation across benchmarks and base models.
+Table 1 presents comprehensive results on two LLMs.
+Across all six settings (3 benchmarks 2 base models), EtCon attains the top Avg.
+score, with large margins over the strongest competing method.
+On Qwen2.5-7B-Instruct, EtCon reaches Avg.
+of 51.5/44.2/56.8 on ZsRE/COUNTERFACT/QAEdit (improving the runner-up by +39.0/+30.9/+35.7).
+On Llama-3-8B-Instruct, EtCon further achieves Avg.
+of 55.6/48.2/55.7 (+35.1/+7.9/+23.2).
+These gains are primarily driven by substantially higher Reliability and Generalization, while maintaining competitive Locality (24.2%–33.6%), indicating that EtCon can internalize new facts and generalize them to paraphrases under sequential edits without overly compromising unrelated behaviors.
+We attribute this robustness to our two-stage design: TPSFT performs localized, stability-aware updates, and GRPO subsequently consolidates the edited knowledge under autoregressive trajectories.
+| Editing Performance | General Capabilities | ||||||||||||
+| Method | Reli. | Gen. | Loc. | Avg. | C-Eval | CoQA | DROP | SQuAD | LogiQA | Avg. | |||
+| Acc. | EM | F1 | EM | F1 | EM | F1 | Acc. | ||||||
+| Pre | 9.8 | 10.1 | 36.2 | 18.7 | 79.5 | 54.5 | 70.1 | 2.2 | 9.9 | 9.9 | 18.9 | 38.7 | 38.0 |
+| FT-M | 14.6 | 14.5 | 30.7 | 19.9 | 75.9 | 21.3 | 38.7 | 2.4 | 13.3 | 2.9 | 11.2 | 37.0 | 25.3 |
+| FT-M+Con. | 42.3 | 34.1 | 31.9 | 36.1 | 77.0 | 26.2 | 46.6 | 2.8 | 14.6 | 4.4 | 13.5 | 37.8 | 27.9 |
+| MMKE | 12.2 | 10.4 | 34.2 | 18.9 | 79.3 | 60.3 | 74.6 | 10.3 | 24.3 | 13.8 | 21.1 | 41.0 | 40.6 |
+| MMKE+Con. | 37.2 | 31.4 | 31.0 | 33.2 | 78.8 | 59.1 | 73.3 | 8.5 | 21.6 | 12.1 | 19.6 | 39.2 | 39.0 |
+| ALPHA | 0.0 | 0.0 | 0.0 | 0.0 | 23.0 | 0.0 | 0.0 | 0.0 | 0.0 | 50.1 | 50.1 | 21.8 | 18.1 |
+| ALPHA+Con. | 0.0 | 0.0 | 0.0 | 0.0 | 23.0 | 0.0 | 0.0 | 0.0 | 0.0 | 50.1 | 50.1 | 21.8 | 18.1 |
+[…]
+| Methods | Edit-wise | 2-hop | 3-hop | 4-hop |
+| Base | - | 69.1 | 36.2 | 85.4 |
+| FT-M | 19.1 | 2.7 | 3.8 | 4.2 |
+| ALPHA | 20.4 | 13.9 | 11.3 | 9.5 |
+| MMKE | 9.3 | 12.0 | 7.1 | 22.9 |
+| MEMIT | 0.4 | 1.1 | 1.3 | 0.0 |
+| EtCon | 54.1 | 24.2 | 14.6 | 33.3 |
+| Stage | Methods | Reli. | Gen. | Loc. | GC-Avg |
+| Base | - | 0.6 | 0.8 | 31.8 | 52.85 |
+| Edit. | w/ SFT | 1.4 | 0.3 | 30.7 | 50.31 |
+| w/ TPSFT | 3.3 | 1.8 | 30.2 | 54.40 | |
+| Con. (after TPSFT) | w/o | 56.1 | 22.4 | 24.7 | 50.17 |
+| w/o | 51.6 | 27.2 | 25.1 | 49.75 | |
+| Complete | 67.1 | 53.4 | 24.2 | 50.25 |
+| Edit Scope | Reli. | Gen. | Loc. | GC-Avg |
+| Full Param. | 47.3 | 41.5 | 35.8 | 58.17 |
+| Target edit | 75.1 | 63.0 | 32.3 | 55.83 |
+Scaling to Longer Editing.
+Figure 5 shows performance evolution when scaling to 3,000 sequential edits on ZsRE.
+EtCon maintains stable performance: Reliability and Generalization decrease moderately from 73.5% to 58.2% and 63.1% to 48.7% respectively, while Locality remains within a narrow band (28%–32%).
+In contrast, FT-M degrades rapidly, with Reliability dropping from 16.6% to near-zero.
+These results demonstrate that EtCon remains stable even under longer editing sequences.
+6 Analysis
+6.1 Effectiveness of Consolidation Stage
+Overall Performance of EtCon.
+As shown in Table 2, EtCon achieves the best trade-off between editing effectiveness and general capability preservation.
+EtCon attains 75.1% Reliability and 63.0% Generalization, substantially outperforming all baselines.
+Meanwhile, EtCon maintains general capabilities close to the original model (C-Eval: 78.5 vs 79.5, CoQA F1: 69.4 vs 70.1), demonstrating that our two-stage design successfully balances knowledge injection with pre-trained capability preservation.
+Performance on Multi-hop Task.
+Table 4 further demonstrates EtCon’s advantage on MQuAKE-CF-v2.
+EtCon achieves 24.2% and 33.3% accuracy on 2-hop and 4-hop queries respectively, substantially outperforming FT-M (2.7%/4.2%).
+This indicates that our consolidation stage can propagate localized edits into more consistent multi-step inference.
+Under the EtCon paradigm, incorporating explicit reward for multi-hop queries during consolidation could further strengthen performance on multi-hop tasks.
+Universality of Knowledge Consolidation.
+To evaluate the generality of our knowledge consolidation stage, we augment three baselines (FT-M, MMKE, and ALPHAEDIT) with Consolidation and compare them with EtCon on QAEdit.
+As shown in Table 2, adding Consolidation to FT-M and MMKE improves Reliability by 25% to 28% and Generalization by about 20%.
+These results suggest that Consolidation bridges the gap between edited parametric knowledge and autoregressive generation behavior, enabling more reliable knowledge utilization in practice.
+Preservation of Pre-trained Capabilities.
+Evaluations across benchmarks show that the knowledge consolidation stage is non-destructive and preserves general capabilities.
+However, it cannot recover capabilities lost during editing: once ALPHAEDIT collapses, the knowledge consolidation stage cannot restore it.
+This indicates that reliable lifelong editing requires both lossless knowledge injection and post-edit consolidation.
+Future work under the EtCon paradigm should prioritize lossless knowledge editing.
+6.2 Ablation Study
+Edit Stage Ablation.
+We compare TPSFT with standard SFT in Table 5.
+The results show that neither method alone yields reliably editable behavior, as evidenced by consistently low Reliability and Generalization.
+Nevertheless, TPSFT offers a clear advantage in capability preservation, substantially reducing the degradation observed with SFT.
+Consolidation Stage Ablation.
+We ablate key components of our consolidation reward (Table 5).
+Removing causes a clear drop (Reliability , Generalization ); inspection shows this encourages “reward hacking,” where the model emits extraneous content (e.g., both old and new facts) to inflate scores.
+Removing is even more detrimental to Reliability (), often producing catastrophic self-contradictions: the model states the correct answer and then immediately reverts.
+These ablations suggest that both reward components help mitigate reward hacking and improve output coherence; see Appendix A.8 for case studies.
+Parameter Sensitivity.
+Figure 4 analyzes TPSFT hyper-parameters.
+Clipping threshold : Increasing improves Reliability and Generalization but reduces Locality, with a sharp drop at .
+Thus, provides the best trade-off, as larger allows stronger updates but increases policy drift.
+Reference update interval: Larger intervals consistently reduce Reliability and Generalization, while Locality remains nearly unchanged.
+Frequent updates (Interval) perform best because stale references increase the policy gap and over-activate clipping, limiting effective updates.
+Effect of CoT Augmentation.
+Figure 6 compares TPSFT with and without CoT-augmented training labels.
+Without CoT augmentation, Reliability is slightly higher, but Locality drops sharply and general capabilities degrade (e.g., lower CoQA compared to base).
+This is because supervising only on the few target tokens leads to overfitting.
+CoT augmentation encourages the model to learn valid reasoning patterns rather than overfitting to isolated input-output mappings, thereby better preserving general capabilities.
+Efficiency of Target Edit.
+We compare Target Edit, which updates a small set of targeted FFN weights, with Full Param update under the same TPSFT setup.
+As shown in Table 6, Target Edit yields substantially higher Reliability and Generalization, while incurring only a slight drop in general capability (GC-Avg).
+In contrast, Full Param attains notably weaker editing performance under the same optimization budget.
+We attribute this gap to reduced optimization efficiency under ratio clipping.
+With limited update steps per edit, Full Param spreads gradients across the whole model, weakening the learning signal on the edited fact.
+In addition, clipping may saturate early for many tokens, further shrinking effective updates.
+Overall, restricting TPSFT provides a more controllable and effective mechanism for knowledge editing.
 7 Conclusion
-In this paper, we identified that the critical gap between theoretical performance and practical effectiveness of knowledge editing methods stems from the absence of a consolidation stage that integrates parametric knowledge into the LLM’s actual generation behavior.
-To address this, we proposed the Edit-then-Consolidate (EtCon) framework, which combines Targeted Proximal Supervised Fine-Tuning (TPSFT) for precise knowledge editing with Group Relative Policy Optimization (GRPO) for effective consolidation.
-TPSFT updates targeted FFN weights within trust-region constraints to ensure reliable edits while preserving pretrained capabilities.
-Then GRPO aligns the edited knowledge with the model’s CoT-based inference policy through a comprehensive reward function.
-Our controlled experiments demonstrated the necessity of the knowledge consolidation stage, and comprehensive evaluations showed that EtCon significantly outperforms existing methods in lifelong editing scenarios, achieving efficient knowledge updates while maintaining locality and preserving model capabilities.
-These results suggest that explicitly decoupling editing and consolidation represents a promising paradigm for practical knowledge editing.
+In this work, we propose EtCon, an edit-then-consolidate framework for reliable knowledge editing.
+In Edit Stage, Targeted Proximal Supervised Fine-Tuning (TPSFT), performs constrained edits to inject knowledge while preserving pre-trained capabilities.
+In the Consolidate Stage, Group Relative Policy Optimization (GRPO), consolidates edits by aligning autoregressive generation with updated facts.
+Extensive experiments demonstrate that EtCon achieves superior editing reliability and generalization while preserving pre-trained capabilities in real-world editing scenarios.
 Appendix A Appendix
 A.1 Additional Implementation Details
+Baseline Configuration.
 For our baseline experiments, we utilize the EasyEdit framework.
-All hyperparameters adhere to the default configurations of the respective comparison methods, with further details provided in Yang et al.
-(2025b); Qi et al.
-(2025).
-For our proposed EtCon method, we update only the FFN down-projection layers (mlp.down_proj) in layers 7–11 of Llama-3-8B-Instruct and layers 5–9 of Qwen2.5-7B-Instruct, following prior works Zhang et al.
-(2024a); Meng et al.
-(2022a); Geva et al.
-(2021); Qi et al.
-(2024) on knowledge editing.
-In the editing stage, we use AdamW with learning rate and set of TPSFT, using an edit batch size of 1 to perform 1,000 sequential single-sample edits on the same model instance.
+All hyperparameters adhere to the default configurations of the respective comparison methods, with further details provided in (Yang et al., 2025b; Qi et al., 2025).
+Editing Stage.
+For our proposed EtCon method, we update only the FFN down-projection layers (mlp.down_proj) in layers 7–11 of Llama-3-8B-Instruct and layers 5–9 of Qwen2.5-7B-Instruct, following prior works (Zhang et al., 2024a; Meng et al., 2022a; Geva et al., 2021; Qi et al., 2024) on knowledge editing.
+Following the lifelong editing paradigm, we use AdamW with learning rate and set of TPSFT, sequentially performing edits (batch size 1) on the same model.
 TPSFT is trained for 5 epochs with at most 6 update steps per edit using early stopping.
-In the consolidation stage, we optimize the inference-time policy with Group Relative Policy Optimization (GRPO).
+Consolidation Stage.
+We optimize the inference-time policy with GRPO.
 The comprehensive reward function in Equation (5) uses the following weight coefficients: for , for , for , and for .
 These weights were determined through extensive empirical experiments to balance factual accuracy with output quality.
+All reward components are binary (0/1), with computed via exact string matching between the extracted answer and the ground-truth after normalization.
 All specific hyperparameters are available in Table 7.
@@
 | Validation Batch Size | 512 |
-| Validation Frequency | Every 5 episodes |
+| Validation Frequency | Every 5 steps |
 | Validation before Training | Yes |
-A.2 Effect of Edited Layers on EtCon
-| Layers | Reli. | Gener. | Local. | |
-Llama-3-8b-Instruct |
+A.2 Additional Results
+A.2.1 Effect of Edited Layers on EtCon
+| Layers | Reli. | Gener. | Local. |
 | 7-8-9-10-11 | 73.5 | 63.1 | 30.2 |
-| 12-13-14-15-16 | 78.1 | 58.3 | 24.1 | |
+| 12-13-14-15-16 | 78.1 | 58.3 | 24.1 |
 | 17-18-19-20-21 | 76.7 | 53.2 | 17.3 |
-We further investigate the impact of editing different FFN layers on EtCon’s performance through an ablation study, as shown in Table 6 and Figure 6.
-We observe that under identical hyperparameter settings, editing early layers (Layers 7-11) outperforms deeper layers (Layers 17-21) in both locality and generalization.
-Analyzing the ”high reward, low performance” phenomenon in deeper layers shown in Figure 6, we find that editing deeper layers is more prone to triggering Reward Hacking.
-Combining existing mechanistic interpretability research—which suggests that shallow layers mainly store factual knowledge while deeper layers handle information integration and reasoning—we posit that this stems from a misalignment between retained knowledge and updated information.
+Experimental Setup. We further investigate the impact of editing different FFN layers on EtCon’s performance through an ablation study, as shown in Table 8 and Figure 7.
+Main Finding. We observe that under identical hyperparameter settings, editing early layers (Layers 7-11) outperforms deeper layers (Layers 17-21) in both locality and generalization.
+Reward Hacking Phenomenon. Analyzing the “high reward, low performance” phenomenon in deeper layers shown in Figure 7, we find that editing deeper layers is more prone to triggering Reward Hacking.
+Mechanistic Interpretation.
+Existing mechanistic interpretability research suggests that shallow layers mainly store factual knowledge while deeper layers handle information integration and reasoning.
+Based on this, we posit that the observed phenomenon stems from a misalignment between retained knowledge and updated information.
+Cognitive Conflict Analysis.
 Editing only deeper layers may cause retained knowledge in shallow layers to conflict with the injected knowledge in deeper layers.
@@
 This reward hacking leads to internal model confusion, thereby causing a decline in performance.
-A.3 Lifelong Editing
-To evaluate the robustness of our framework in larger-scale lifelong editing scenarios, we extend the ZsRE sequential editing experiments to 3,000 single-sample edits on distinct ZsRE examples.
-At each checkpoint (600, 1200, 1800, 2400, and 3000 edits), we first run the knowledge consolidation phase and then evaluate the edited LLM before proceeding with further edits, as illustrated in Figure 7.
-Across all three metrics, EtCon exhibits graceful degradation: starting from the first 600 edits, its Reliability and Generalization remain high and decrease only moderately as the stream length triples, while Locality fluctuates within a narrow band without signs of collapse.
-In contrast, FT-M begins with much lower scores and deteriorates rapidly as more edits accumulate, with Reliability and Generalization approaching near-zero and Locality dropping sharply.
-These results indicate that EtCon can sustain thousands of sequential edits and still substantially outperform the baseline under lifelong editing scenarios.
-A.4 Evaluation on Reasoning-Oriented Architectures
-| Layers | Reli. | Gener. | Local. | |
-Deepseek-R1-distill-Qwen-7B
+A.2.2 Portability Evaluation
+Superior Cross-Architecture Portability.
+As shown in the Figure 8, EtCon demonstrates superior portability across architectures, achieving 26.2% on Llama-3 and 24.3% on Qwen2.5, consistently outperforming the strongest baseline FT-M (18.8% and 21.8%).
+Consolidation Effectiveness. This validates that the consolidation stage effectively enables the model to utilize edited facts for reasoning.
+Limitations and Future Directions.
+However, the absolute gains remain limited compared to single-hop reliability.
+We attribute this to EtCon being trained solely on single-hop QA without explicit supervision for multi-hop compositionality; thus, incorporating such supervision is a natural next step to further enhance portability.
+A.3 Evaluation on Reasoning-Oriented Architectures
+| Layers | Reli. | Gener. | Local. |
 | 5-6-7-8-9 | 88.6 | 53.5 | 17.0 |
-| 13-14-15-16-17 | 79.6 | 42.6 | 16.1 | |
+| 13-14-15-16-17 | 79.6 | 42.6 | 16.1 |
 | 23-24-25-26-27 | 83.0 | 52.9 | 7.3 |
+Experimental Setup.
 To assess the robustness of our method on architectures with inherent reasoning capabilities, we extended our evaluation to DeepSeek-R1-Distill-Qwen-7B using 1,000 samples from the ZsRE.
-As reported in Table 7, our method maintains high efficacy on this reasoning-oriented architecture.
+Following our main experiments, we target the MLP down-projection weights (mlp.down_proj) for knowledge editing.
+Shallow Layers Yield Best Trade-off.
 As reported in Table 9, EtCon maintains high editing efficacy on this reasoning-oriented model.
@@
 This suggests that our Edit-then-Consolidate paradigm is compatible with the model’s intrinsic reasoning processes rather than disrupting them.
+Deep-Layer Shortcut Phenomenon.
 The green curve exhibits the fastest initial convergence (Steps 0-20).
 Since these layers are proximal to the output projection, the model can quickly maximize the reward by establishing a direct mapping to the target answer.
-However, similar to the phenomenon observed in Appendix A.2, this configuration plateaus at a lower reliability level.
+However, similar to the phenomenon observed in Appendix A.2.1, this configuration plateaus at a lower reliability level.
 This suggests that modifying only the deep layers could lead the model to adopt a ”shortcut” strategy which may leave the model vulnerable to the cognitive conflict between retained shallow knowledge and updated deep injections.
+Shallow-Layer Constructive Consolidation.
 In contrast, the red curve displays a distinct ”warm-up” phase (Steps 0-15) followed by a sustained ascent to the highest performance plateau.
@@
 This results in convergence stability and reliability, demonstrating that our Edit-then-Consolidate paradigm is compatible with the intrinsic reasoning mechanisms of complex architectures.
-A.5 Time Efficiency Analysis
+A.4 Time Efficiency Analysis
 | Methods | TPSFT (Ours) | AlphaEdit | MEMIT | GRACE | WISE | FT-M |
@@
 These results indicate that, under the same configuration, TPSFT+Con attains faster convergence in the consolidation stage than the baselines.
-A.6 CoT Generation and Processing
+A.5 CoT Generation and Processing
 To generate Chain-of-Thought reasoning paths during TPSFT training, we employ the following prompt template that encourages natural reasoning while ensuring convergence to target answers:
@@
 We additionally discard CoT samples whose final answer is clearly inconsistent with the ground-truth and regenerate them, further reducing the risk of noisy supervision.
-A.7 Real-world evaluation Details
-In this work, we follow the design Yang et al.
-(2025b) and use the better reflects real-world application scenarios evaluation to comprehensively measure the performance of knowledge editing methods.
+A.6 Real-world evaluation Details
+In this work, we follow the design (Yang et al., 2025b) and use the better reflects real-world application scenarios evaluation to comprehensively measure the performance of knowledge editing methods.
 Specifically, our evaluation process consists of three key stages:
@@
 (3) LLM-as-a-Judge Evaluation: To achieve a scalable and objective evaluation, we introduce a more powerful Large Language Model (LLM) to act as a ”judge.” This judge model makes its decision by comprehensively considering the original question, the ground-truth answer (Target), and the full answer content from the edited model, ultimately providing a binary (correct/incorrect) judgment.
-The full judge prompts are provided in Fig 4 and Fig.
-5
+The full judge prompts are provided in Figure 11 and Figure 12
+A.7 Metrics for MQuAKE-CF-v2
+For multi-hop task evaluation, we sample 1000 instances from MQuAKE-CF-v2 proportionally across 2-hop, 3-hop, and 4-hop questions. We report the following metrics:
+Edit-wise accuracy: measures whether the model correctly answers direct queries about edited facts.
+Multi-hop accuracy (2/3/4-hop): evaluates whether the model can perform multi-step reasoning that involves the edited knowledge.
 A.8 Analysis of Reward Hacking Cases
 Analysis of Reward Hacking Patterns: The two cases in Figures.
-10 and 11 reveal distinct failure modes in the absence of proper reward design.
+13 and 14 reveal distinct failure modes in the absence of proper reward design.
 In Figure 13, the model exhibits ”self-correction” behavior—correctly reasoning through the problem but then artificially inserting the target answer followed by an immediate correction.